### v1

# Entire Large Solutions for Competitive Semilinear Elliptic Systems with General Nonlinearities Satisfying Keller–Osserman Conditions

Abstract
We generalize a theorem of Lair concerning the existence of positive entire large solutions to competitive semilinear elliptic systems. While Lair’s original result [3] was established for power-type nonlinearities, our work extends the theory to a broad class of general nonlinearities satisfying a Keller–Osserman-type growth condition. The proof follows the same conceptual framework monotone iteration to construct global positive solutions, reduction to a scalar inequality for the sum of the components, application of a Keller–Osserman transform, and a two-step radial integration argument but replaces the explicit power-law growth with a general monotone envelope function. This approach yields a unified and verifiable criterion for the existence of large solutions in terms of the Keller–Osserman integral, thereby encompassing both critical and supercritical growth regimes within a single analytical setting.
MSC 2020 Subject Classification: 35J47, 35J57, 35B40, 35B44.
Keywords and Phrases: Semilinear elliptic systems; entire large solutions; Keller–Osserman condition; monotone iteration; radial solutions; blow-up behaviour; competitive systems; nonlinear growth conditions.
1 Introduction
We investigate the existence of positive entire large solutions for the radial semilinear elliptic system
$\begin{cases}\Delta u=p(r)f(u,v),\\ \Delta v=q(r)g(u,v),\end{cases}\quad r=|x|,\quad x\in\mathbb{R}^{n},\ n\geq 3,$ (1.1)
under the following assumptions:
  • (H1)
    (Weights) The functions $p,q:[0,\infty)\rightarrow[0,\infty)$ are continuous, not both identically zero. We define the Green-type radial potentials
    $P(\infty)=\frac{1}{n-2}\int_{0}^{\infty}r\,p(r)\,dr<\infty,\quad Q(\infty)=\frac{1}{n-2}\int_{0}^{\infty}r\,q(r)\,dr<\infty,$
    where
    $P(r):=\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,ds\,dt,\quad Q(r):=\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,ds\,dt,$
    and $P(\infty)=\lim_{r\rightarrow\infty}P(r)$, $Q(\infty)=\lim_{r\rightarrow\infty}Q(r)$. We assume $\min(p,q)$ is not compactly supported.
  • (H2)
    (Monotonicity) The nonlinearities $f,g:[0,\infty)\times[0,\infty)\rightarrow[0,\infty)$ are nondecreasing in each variable, satisfy $f\left(s,t\right)>0$ and $g\left(s,t\right)>0$ $\forall s,t\in\left(0,\infty\right)$, $f(0,0)=g(0,0)=0$, are continuous on $0,\infty)^{2}$ and differentiable on $(0,\infty)^{2}$.
  • (H3)
    (Boundedness and Keller–Osserman envelope) There exists a continuous, nondecreasing function $h:[0,\infty)\rightarrow[0,\infty)$ with $h\left(s\right)>0$ $\forall s\in\left(0,\infty\right)$, $h\left(0\right)=0$ and constants $\eta>0$ sufficiently small such that
    $f(s,t)+g(s,t)\geq h(s+t),\quad\forall s,t\geq\eta.$
    Moreover, the Keller–Osserman integral
    $H\left(\infty\right)=\lim_{s\rightarrow\infty}\int_{1}^{s}\frac{dt}{\sqrt{\int_{0}^{t}h(z)\,dz}}<\infty,\qquad$ (1.2)
    is finite.
Our aim is to extend the existence results of Lair [3] originally established for power-type nonlinearities to a broad class of general nonlinearities satisfying a Keller–Osserman-type growth condition. This generalization unifies the treatment of critical and supercritical growth regimes and provides a framework that encompasses earlier results such as those of Cîrstea and Rădulescu [1], Keller [2], Osserman [5], and others.
A positive entire solution of (1.1) is a pair $(u,v)$ of $C^{2}$ radial functions such that $u(r)>0$ and $v(r)>0$ for all $r\geq 0$. We denote $\mathbb{R}^{+}=[0,\infty)$ and define the set of central values
[…]
The structure of $\mathcal{G}$ plays a central role in our analysis. In particular, we establish monotonicity properties, boundedness, and closedness of certain subsets of $\mathcal{G}$, and we identify a boundary set where all corresponding solutions are large, i.e., at least one component diverges to infinity as $r\rightarrow\infty$.
Our approach adapts the barrier method and radial integration techniques of Lair [4, 3], replacing explicit power-law growth with a general monotone envelope function $h$ (and implicitly $f+g$) satisfying the Keller–Osserman condition. This allows us to recover, as special cases, known results for power-type nonlinearities [3] and to connect with complementary bounded-solution results such as those in [8, 6, 7].
Paper structure. The paper is organized as follows: Section 2 presents our main results. Section 3 contains the detailed proofs of these results. The paper ends with Section 4, where we present our concluding remarks.
2 Main Results
We begin with the following property of the set $\mathcal{G}$.
Theorem 2.1.
Let $\alpha_{0}>0$, $\beta_{0}>0$. Then $\mathcal{G}$ is nonempty and has the property that if $(\alpha_{0},\beta_{0})\in\mathcal{G}$ and $0<\alpha<\alpha_{0}$, $0<\beta<\beta_{0}$, then $(\alpha,\beta)\in\mathcal{G}$.
For $\delta>0$ sufficiently small, define
$\mathcal{G}_{\delta}:=\left\{(\alpha,\beta)\in\mathcal{G}\ \middle|\ \min\{\alpha,\beta\}\geq\delta\right\}.$
We have the following result.
Theorem 2.2.
The set $\mathcal{G}_{\delta}$ is both closed and bounded in $(0,\infty)\times(0,\infty)$.
Next, we define the edge set
$\mathcal{E}(\mathcal{G}_{\delta}):=\left\{(\alpha,\beta)\in\partial\mathcal{G}_{\delta}\ \middle|\ \min\{\alpha,\beta\}>\delta\right\}.$
We can now state our main existence theorem for entire large solutions of (1.1).
Theorem 2.3.
There exists a positive radial solution $(u,v)$ of (1.1) with central value $(u(0),v(0))\in\mathcal{E}(\mathcal{G}_{\delta})$ such that at least one of $u$ or $v$ diverges, i.e.,
$\lim_{r\rightarrow\infty}u(r)=\infty\quad\text{or}\quad\lim_{r\rightarrow\infty}v(r)=\infty.$ (2.1)
Moreover, if
$H(s):=\int_{s}^{\infty}\frac{dt}{f(t,t)+g\left(t,t\right)}$
and $H^{-1}$ denotes its inverse, then for all sufficiently large $r$,
$u(r)+v(r)\ \geq\ H^{-1}\!\left(\int_{r}^{\infty}t^{1-n}\int_{0}^{t}s^{n-1}\big{(}p(s)+q(s)\big{)}\,ds\,dt\right).$ (2.2)
If, in addition, we impose the following Keller-Osserman condition at infinity:
[…]
then
$\lim_{r\rightarrow\infty}u(r)=\lim_{r\rightarrow\infty}v(r)=\infty.$
Remark 2.4 (Relation to Lair’s result).
When $f(u,v)=u^{a}v^{b}$ and $g(u,v)=u^{c}v^{d}$ with $\min\{a,d\}>1$, the envelope $h(s)=s^{\min\{a+b,c+d\}}$ satisfies the Keller–Osserman condition if and only if $\min\{a+b,c+d\}>1$. This recovers the main theorem of [3] as a special case.
Remark 2.5 (Relation to Zhang et al.’s result).
Both Theorem 2.3 and Theorem 1.2 of [8] study radial elliptic systems of the form (1.1). The two results are therefore compatible: they describe different behaviours (bounded vs. large) in disjoint regions of the central value space, even for the same coefficients.
In Theorem 1.2 of [8], the focus is on positive bounded entire radial solutions, under a strict hierarchy between the finite numbers $P(\infty)$, $Q(\infty)$ and $H(\infty)$.
Theorem 2.3 asserts that for $(u(0),v(0))\in\mathcal{E}(\mathcal{G}_{\delta})$, any entire solution is large, while Theorem 1.2 of [8] produces bounded entire solutions for $(u(0),v(0))$ in the interior of $\mathcal{G}$.
3 Proof of the Theorems
We begin by recalling that any radial $C^{2}$ solution $(u,v)$ of (1.1) satisfies the equivalent system of Volterra-type integral equations
$u(r)$ $=\alpha_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,f\big{(}u(s),v(s)\big{)}\,ds\,dt,$ (3.1)
$v(r)$ $=\beta_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,g\big{(}u(s),v(s)\big{)}\,ds\,dt,$ (3.2)
valid for all $r\geq 0$. Here $\alpha_{0}=u(0)$ and $\beta_{0}=v(0)$ are the central values. It is clear that for $\alpha_{0},\beta_{0}\in(0,\infty)$, any solution of (3.1)–(3.2) is also a solution of (1.1), although it may not necessarily satisfy the large-solution condition (2.1).
Proof of Theorem 2.1
Proof.
The argument follows the monotone iteration scheme of [3], adapted to the present setting where $f$ and $g$ satisfy the general hypotheses (H2)–(H3) rather than being pure power functions.
Step 1: Construction of the iterative sequences. Given $(\alpha_{0},\beta_{0})\in(0,\infty)\times(0,\infty)$, define
$u_{0}(r)$ $\equiv\alpha_{0}>0,$ $v_{0}(r)$ $\equiv\beta_{0}>0,$
$u_{k}(r)$ $=\alpha_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,f\big{(}u_{k-1}(s),v_{k-1}(s)\big{)}\,ds\,dt,$
$v_{k}(r)$ $=\beta_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,g\big{(}u_{k-1}(s),v_{k-1}(s)\big{)}\,ds\,dt,$
for $k\geq 1$.
Step 2: Monotonicity. By the positivity of $p,q$ and the monotonicity of $f,g$ in each variable (assumption (H2)), we have
$u_{k}(r)\geq u_{k-1}(r),\quad v_{k}(r)\geq v_{k-1}(r),\quad\forall r\geq 0,\ \forall k\geq 1.$
Thus, for each fixed $r$, both sequences $\{u_{k}(r)\}$ and $\{v_{k}(r)\}$ are nondecreasing.
Step 3: Uniform upper bound via a scalar supersolution. The Keller–Osserman condition (H2) ensures the existence of a scalar function $w\in C^{2}([0,\infty))$ satisfying
$\Delta w=(p+q)\big{(}f(w,w)+g(w,w)\big{)}\quad\text{in }[0,\infty),\quad w(0)>\alpha_{0}+\beta_{0},\quad\lim_{r\to\infty}w(r)=\infty.$ (3.3)
Such a $w$ can be constructed by solving the associated scalar radial problem with initial value $w(0)$ large enough and using the Keller–Osserman integral to guarantee blow-up at infinity.
By the comparison principle for radial integral equations, we have
$u_{k}(r)+v_{k}(r)\leq w(r),\quad\forall r\geq 0,\ \forall k\geq 0.$
This provides a uniform bound on the iterates on every compact interval.
Step 4: Convergence. Since the sequences are monotone increasing and uniformly bounded above on compact intervals, they converge pointwise to limits
$u(r):=\lim_{k\rightarrow\infty}u_{k}(r),\quad v(r):=\lim_{k\rightarrow\infty}v_{k}(r),\quad r\geq 0.$
By monotone convergence, the limits $(u,v)$ satisfy (3.1)–(3.2) for all $r\geq 0$.
Step 5: Conclusion. The limit pair $(u,v)$ is a positive $C^{2}$ radial solution of (1.1) with $(u(0),v(0))=(\alpha_{0},\beta_{0})$. Thus $(\alpha_{0},\beta_{0})\in\mathcal{G}$, proving that $\mathcal{G}$ is nonempty.
Step 6: Downward closure property. If $(\alpha_{0},\beta_{0})\in\mathcal{G}$ and $0<\alpha<\alpha_{0}$, $0<\beta<\beta_{0}$, then starting the iteration from $(\alpha,\beta)$ yields smaller initial data. By the same monotonicity and boundedness arguments, the iteration converges to a positive entire solution, hence $(\alpha,\beta)\in\mathcal{G}$.
This completes the proof of Theorem 2.1. ∎
3.1 Proof of Theorem 2.2
Proof.
We divide the proof into two parts: boundedness and closedness of $\mathcal{G}_{\delta}$.
Boundedness. Suppose, for contradiction, that $\mathcal{G}_{\delta}$ is unbounded in $(0,\infty)\times(0,\infty)$. From Theorem 2.1, we know that $\mathcal{G}$ is downward closed in each coordinate: if $(\alpha_{0},\beta_{0})\in\mathcal{G}$ and $0<\alpha<\alpha_{0}$, $0<\beta<\beta_{0}$, then $(\alpha,\beta)\in\mathcal{G}$.
In particular, if $(\alpha_{0},\beta_{0})\in\mathcal{G}_{\delta}$, then for any $(\alpha,\beta)$ with
$\delta\leq\alpha\leq\alpha_{0},\quad\delta\leq\beta\leq\beta_{0},$
we also have $(\alpha,\beta)\in\mathcal{G}_{\delta}$. This implies that if $\mathcal{G}_{\delta}$ were unbounded, it would necessarily contain an unbounded subset of the form
$\delta,\infty)\times\{\delta\}\quad\text{or}\quad\{\delta\}\times[\delta,\infty).$
Without loss of generality, assume that
$\{\delta\}\times[\delta,\infty)\subseteq\mathcal{G}_{\delta}.$
Step 1: Choice of a comparison function. Let $\widetilde{p}(r):=\min\{p(r),q(r)\}$. By assumption (H1), $\widetilde{p}$ is continuous, radial, and not compactly supported. Thus, there exists $R>0$ such that $\widetilde{p}(R)>0$.
From [4], there exists a large, nonnegative radial solution $w$ of
$\Delta w=\widetilde{p}(r)\,h(w)\quad\text{in }[0,R),\quad w(0)=\delta,\quad\lim_{r\rightarrow R^{-}}w(r)=\infty.$
Here “large” means that $w$ blows up at the finite radius $R$.
Step 2: Construction of a solution in $\mathcal{G}_{\delta}$. Since $\{\delta\}\times[\delta,\infty)\subseteq\mathcal{G}_{\delta}$, pick $\alpha_{0}>\delta$ and consider the initial data
$u(0)=\alpha_{0}>w(0)=\delta,\quad v(0)=\delta.$
By Theorem 2.1, there exists a positive entire solution $(u,v)$ of (3.1)–(3.2) with these initial values, i.e.,
$u(r)$ $=\alpha_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,f\big{(}u(s),v(s)\big{)}\,ds\,dt,\quad 0\leq r<\infty,$
$v(r)$ $=\delta+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,g\big{(}u(s),v(s)\big{)}\,ds\,dt,\quad 0\leq r<\infty.$
Step 3: Differential inequality for $u+v$. Since $u(r)\geq\alpha_{0}>\delta$ and $v(r)\geq\delta$ for all $r\geq 0$, we have
$u(r)+v(r)\geq\alpha_{0}+\delta>2\delta.$
By assumption (H3), $f(s,t)+g(s,t)\geq h(s+t)$ for $s,t\geq\eta$ (and here $\eta\leq\delta$), so
$\Delta(u+v)=p(r)f(u,v)+q(r)g(u,v)\geq\widetilde{p}(r)\,h(u+v)\quad\text{in }[0,R).$
Step 4: Comparison with $w$. The function $w$ satisfies $\Delta w=\widetilde{p}(r)\,h(w)$ in $[0,R)$, $w(0)=\delta$, and blows up at $r=R$. By the strong maximum principle (applied to u+vwu+v-w), if $u(0)+v(0)>w(0)$, then $u+v>w$ in a neighborhood of $r=0$. However, since $w$ blows up at $R$ while $u+v$ is $C^{2}$ on $[0,R)$, the comparison principle forces $u+v\leq w$ in $[0,R)$. This is a contradiction because
$u(0)+v(0)=\alpha_{0}+\delta>\delta=w(0).$
Therefore, our assumption that $\mathcal{G}_{\delta}$ is unbounded must be false, and $\mathcal{G}_{\delta}$ is bounded.
Closedness. Let $\{(\alpha_{m},\beta_{m})\}\subset\mathcal{G}_{\delta}$ be a sequence converging to $(\alpha,\beta)\in\overline{\mathcal{G}_{\delta}}$. For each $m$, Theorem 2.1 guarantees the existence of a positive entire solution $(u_{m},v_{m})$ of (3.1)–(3.2) with
$u_{m}(0)=\alpha_{m},\quad v_{m}(0)=\beta_{m}.$
From the monotone iteration scheme in Theorem 2.1 and the Keller–Osserman supersolution $w$ from (3.3) with $w(0)>\alpha_{m}+\beta_{m}$, we deduce that the family $\{(u_{m},v_{m})\}$ is uniformly bounded on compact intervals and equicontinuous. By the Arzelá –Ascoli theorem, there exists a subsequence converging uniformly on compact sets to a limit $(u,v)$.
Passing to the limit in (3.1)–(3.2) shows that $(u,v)$ is a positive entire solution with
$u(0)=\alpha,\quad v(0)=\beta.$
Since $\min\{\alpha,\beta\}\geq\delta$ by construction, we have $(\alpha,\beta)\in\mathcal{G}_{\delta}$. Thus $\mathcal{G}_{\delta}$ is closed.
This completes the proof of Theorem 2.2. ∎
3.2 Proof of Theorem 2.3
Proof.
Standing assumptions and notation. We work under hypotheses (H1)–(H3) and consider the radial system (1.1) in its equivalent integral (Green) form (3.1)–(3.2), with prescribed central values
$(u(0),v(0))=(\alpha,\beta)\in\mathcal{E}(\mathcal{G}_{\delta}).$
Define the Keller–Osserman transform
$H(s):=\int_{s}^{\infty}\frac{dt}{f(t,t)+g\left(t,t\right)},\quad s>0.$
By construction, $H$ is strictly decreasing,
$H^{\prime}(s)=\frac{-1}{f(s,s)+g\left(s,s\right)}<0,$
and $H^{\prime\prime}(s)\geq 0$ since $s\rightarrow f(s,s)+g\left(s,s\right)$ is increasing.
Step 1: Perturbation of initial data and blow-up for approximants. From Theorem 2.1, $\mathcal{G}$ is nonempty and downward closed in each coordinate. From Theorem 2.2, $\mathcal{G}_{\delta}$ is closed and bounded for $\delta>0$ sufficiently small.
Take $(\alpha,\beta)\in\mathcal{E}(\mathcal{G}_{\delta})$. For each $k\geq 1$, consider the perturbed initial data
$\left(\alpha+\frac{1}{k},\ \beta+\frac{1}{k}\right).$
If this pair belonged to $\mathcal{G}$, then by downward closure and the definition of $\mathcal{G}_{\delta}$, $(\alpha,\beta)$ would lie in the interior of $\mathcal{G}_{\delta}$, contradicting $(\alpha,\beta)\in\mathcal{E}(\mathcal{G}_{\delta})$. Hence $(\alpha+\frac{1}{k},\beta+\frac{1}{k})\notin\mathcal{G}$.
Therefore, the solution $(u_{k},v_{k})$ of
$u_{k}(r)$ $=\alpha+\frac{1}{k}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)f(u_{k}(s),v_{k}(s))\,ds\,dt,$
$v_{k}(r)$ $=\beta+\frac{1}{k}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)g(u_{k}(s),v_{k}(s))\,ds\,dt,$
must blow up in finite radius: there exists $R_{k}>0$ such that
$\lim_{r\rightarrow R_{k}^{-}}u_{k}(r)=\infty\quad\text{or}\quad\lim_{r\rightarrow R_{k}^{-}}v_{k}(r)=\infty.$
In either case,
$\lim_{r\rightarrow R_{k}^{-}}\big{(}u_{k}(r)+v_{k}(r)\big{)}=\infty.$
Moreover, by monotonicity in $k$ of the blow-up radii (see [1]),
$R_{1}\leq R_{2}\leq\dots\leq R_{k}\leq\dots,$
and we set
$R:=\lim_{k\rightarrow\infty}R_{k}\in[R_{1},\infty].$
Step 2: Reduction to a scalar inequality. Let
$W_{k}(r):=u_{k}(r)+v_{k}(r),\quad W(r):=u(r)+v(r),$
where $(u,v)$ is the limit solution corresponding to $(\alpha,\beta)$. Since $u_{k}\leq W_{k}$ and $v_{k}\leq W_{k}$, and $f,g$ are nondecreasing,
$\Delta W_{k}$ $=p(r)f(u_{k},v_{k})+q(r)g(u_{k},v_{k})$
$\leq(p(r)+q(r))\,\left(f(W_{k},W_{k})+g\left(W_{k},W_{k}\right)\right),$ (3.4)
where we used (H2) in the last step.
Step 3: Applying the Keller–Osserman transform. Applying $H$ to $W_{k}$ and using $H^{\prime\prime}\geq 0$, the chain rule gives
$\Delta H(W_{k})=H^{\prime}(W_{k})\,\Delta W_{k}+H^{\prime\prime}(W_{k})|\nabla W_{k}|^{2}\geq H^{\prime}(W_{k})\,\Delta W_{k}.$
Since $H^{\prime}(s)=-1/\left(f(s,s)+g\left(s,s\right)\right)$, inequality (3.4) yields
$\Delta H(W_{k})\geq-(p(r)+q(r)).$
Step 4: First radial integration. In radial coordinates, the above inequality becomes
$\big{(}r^{n-1}(H(W_{k}))^{\prime}(r)\big{)}^{\prime}\geq-r^{n-1}(p(r)+q(r)).$
Integrating from 0 to $r<R_{k}$ and using regularity at $r=0$ (which implies $\lim_{t\downarrow 0}t^{n-1}(H(W_{k}))^{\prime}(t)=0$), we obtain
$(H(W_{k}))^{\prime}\geq-r^{1-n}\int_{0}^{r}s^{n-1}(p(s)+q(s))\,ds.$
Step 5: Second radial integration and lower bound. Integrating from $r$ to $R_{k}$ gives
$H(W_{k}(R_{k}))-H(W_{k}(r))\geq-\int_{r}^{R_{k}}t^{1-n}\left(\int_{0}^{t}s^{n-1}(p(s)+q(s))\,ds\right)dt.$
Letting $r\rightarrow R_{k}^{-}$, we have $W_{k}(R_{k})\rightarrow\infty$ and hence $H(W_{k}(R_{k}))\rightarrow 0$. Thus
$H(W_{k}(r))\leq\int_{r}^{R_{k}}t^{1-n}\left(\int_{0}^{t}s^{n-1}(p(s)+q(s))\,ds\right)dt.$
Since $H$ is strictly decreasing, we invert to obtain
$W_{k}(r)\geq H^{-1}\!\left(\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}(p(s)+q(s))\,ds\,dt\right).$
Passing to the limit $k\rightarrow\infty$ yields
$W(r)\geq H^{-1}\!\left(\int_{r}^{R}t^{1-n}\int_{0}^{t}s^{n-1}(p(s)+q(s))\,ds\,dt\right).$
If $R<\infty$, the right-hand side tends to $\infty$ as $r\rightarrow R^{-}$, contradicting the fact that $(u,v)$ is entire. Hence $R=\infty$, and we have established the quantitative lower bound (2.2).
Step 6: Divergence of at least one component. Proceeding as in Step 1 of the proof of Theorem 2.2, we have already shown that
$W(r):=u(r)+v(r)\;\longrightarrow\;+\infty\quad\text{as }r\rightarrow\infty.$
If both $u$ and $v$ were bounded on $0,\infty)$, say $u(r)\leq M_{1}$ and $v(r)\leq M_{2}$ for all $r\geq 0$, then
$W(r)\leq M_{1}+M_{2}\quad\forall r\geq 0,$
contradicting $W(r)\rightarrow+\infty$. Hence at least one of $u,v$ must diverge:
$\lim_{r\rightarrow\infty}u(r)=+\infty\quad\text{or}\quad\lim_{r\rightarrow\infty}v(r)=+\infty.$
Both components diverge under the supplementary Keller–Osserman condition. Assume now the additional Keller–Osserman condition at infinity (2.3). We prove
$\lim_{r\rightarrow\infty}u(r)\;=\;\lim_{r\rightarrow\infty}v(r)\;=\;+\infty.$
Clearly, if $u$ were bounded, then $v=W-u\rightarrow+\infty$. Thus it suffices to show
$v(r)\rightarrow+\infty\ \Longrightarrow\ u(r)\rightarrow+\infty,\qquad u(r)\rightarrow+\infty\ \Longrightarrow\ v(r)\rightarrow+\infty.$
Claim 1: If $v(r)\rightarrow+\infty$ as $r\rightarrow\infty$, then $u(r)\rightarrow+\infty$ as $r\rightarrow\infty$.
Proof. Assume by contradiction that $u$ is bounded: there exists $U>0$ with $0<u(r)\leq U$ for all $r\geq 0$. Since $v(r)\rightarrow+\infty$ while $u$ is bounded, there exists $R_{0}>0$ such that
$u(r)\leq U<v(r)\qquad\forall r\geq R_{0}.$
By the monotonicity of $f$ in each variable,
$f\bigl{(}u(r),v(r)\bigr{)}>f\bigl{(}u(r),u(r)\bigr{)}=:\ell_{f}\bigl{(}u(r)\bigr{)}\qquad\forall r\geq R_{0},$
where $\ell_{f}(s):=f(s,s)$ is continuous, nondecreasing, $\ell_{f}(0)=0$, and satisfies the Keller–Osserman condition
$\int_{1}^{\infty}\frac{dt}{\sqrt{\int_{0}^{t}\ell_{f}(z)\,dz}}<\infty.$
Thus, for $r\geq R_{0}$,
$\Delta u(r)>p(r)\,\ell_{f}\bigl{(}u(r)\bigr{)}.$ (3.5)
We now carry out a Keller–Osserman barrier construction with smaller initial data. Fix $R_{0}>0$ as above and consider the radial Cauchy problem
$y^{\prime\prime}(r)+\dfrac{n-1}{r}y^{\prime}(r)$ $=$ $p(r)\,\ell_{f}\!\big{(}y(r)\big{)},\text{ in }\left[R_{0},R_{a}\right)$ (3.6)
$y(R_{0})$ $=$ $a<u(R_{0}),\quad y^{\prime}(R_{0})=0,$
which has a (maximal) $C^{2}$ solution $y$ on $\left[R_{0},R_{a}\right)$ by standard ODE theory.
Define the Keller–Osserman primitive and we denote
$\mathcal{C}_{f}(s):=\int_{s_{\ast}}^{s}\frac{d\tau}{\ell_{f}(\tau)},\qquad s>s_{\ast},$
where $s_{\ast}>0$ is chosen so that $\ell_{f}(\tau)>0$ on $s_{\ast},\infty)$. Then $\mathcal{C}_{f}$ is strictly increasing on $(s_{\ast},\infty)$, with
$\mathcal{C}_{f}^{\prime}(s)=\frac{1}{\ell_{f}(s)},\qquad\mathcal{C}_{f}^{\prime\prime}(s)=-\frac{\ell_{f}^{\prime}(s)}{\ell_{f}(s)^{2}}\leq 0,$
and, by the Keller–Osserman condition, $\mathcal{C}_{f}(\infty):=\lim_{s\rightarrow\infty}\mathcal{C}_{f}(s)<\infty$.
Set $W(r):=\mathcal{C}_{f}\bigl{(}y(r)\bigr{)}$. Differentiating and using (3.6) yields
$W^{\prime}(r)=\frac{y^{\prime}(r)}{\ell_{f}(y(r))},\qquad\bigl{(}r^{n-1}W^{\prime}(r)\bigr{)}^{\prime}=\frac{r^{n-1}}{\ell_{f}(y)}\Bigl{(}y^{\prime\prime}+\frac{n-1}{r}y^{\prime}\Bigr{)}-\frac{r^{n-1}\ell_{f}^{\prime}(y)}{\ell_{f}(y)^{2}}\,y^{\prime 2}\;\leq\;r^{n-1}p(r).$
Integrating from $R_{0}$ to $r$ and using $W^{\prime}(R_{0})=y^{\prime}(R_{0})/\ell_{f}(a)=0$ gives
$r^{n-1}W^{\prime}(r)\;\leq\;\int_{R_{0}}^{r}s^{n-1}p(s)\,ds.$
Dividing by $r^{n-1}$ and integrating once more from $R_{0}$ to $r$ leads to the Keller–Osserman estimate
$\mathcal{C}_{f}\bigl{(}y(r)\bigr{)}\;\leq\;\mathcal{C}_{f}(a)\;+\;Q_{R_{0}}(r),\qquad Q_{R_{0}}(r):=\int_{R_{0}}^{r}t^{1-n}\!\left(\int_{R_{0}}^{t}s^{n-1}p(s)\,ds\right)\,dt.$ (3.7)
By direct computation (Fubini’s theorem and $n\geq 3$),
$0\leq Q_{R_{0}}(\infty)\;=\frac{1}{n-2}\int_{R_{0}}^{\infty}\left(1-\left(\frac{R_{0}}{t}\right)^{n-2}\right)t\,p(t)\,dt\;\leq\;\frac{1}{n-2}\int_{R_{0}}^{\infty}t\,p(t)\,dt\;,$
so $Q_{R_{0}}(\infty)<\infty$, and moreover $Q_{R_{0}}(\infty)\downarrow 0$ as $R_{0}\uparrow\infty$.
Choosing $R_{0}$ large so that $Q_{R_{0}}(\infty)$ is small, and $a$ close enough to $u(R_{0})$ but still smaller than $u(R_{0})$, we can ensure
$\mathcal{C}_{f}(a)+Q_{R_{0}}(\infty)>\mathcal{C}_{f}(\infty),$
which forces blow-up of $y$ at some finite $R_{a}>R_{0}$.
Now compare $u$ and $y$ on $A_{R_{0},R}$ for $R<R_{a}$. On $\partial A_{R_{0},R}$ we have $u(R_{0})>a=y(R_{0})$ and, for $R$ close to $R_{a}$, $y(R)$ arbitrarily large while $u$ is bounded. By the weak maximum principle applied to uyu-y in $\{u<y\}$ (note the reversed inequality at $R_{0}$), we reach a contradiction with the boundedness of $u$ (as in Step 1 of the proof of Theorem 2.2). Therefore $u$ must be unbounded, hence $u(r)\rightarrow+\infty$ as $r\rightarrow\infty$. ∎
Claim 2: If $u(r)\rightarrow+\infty$ as $r\rightarrow\infty$, then $v(r)\rightarrow+\infty$ as $r\rightarrow\infty$.
Proof. The argument is symmetric, using $\ell_{g}$, $q$, and a barrier $w$ with $w(R_{1})<v(R_{1})$ to obtain the same type of contradiction. ∎
Conclusion. Combining the two claims yields, under (2.3),
$\lim_{r\rightarrow\infty}u(r)\;=\;\lim_{r\rightarrow\infty}v(r)\;=\;+\infty.$
Step 7: Conclusion. We have constructed a positive radial entire solution $(u,v)$ of (1.1), established
$W\left(r\right)=u\left(r\right)+v\left(r\right)\rightarrow\infty\text{ as }r\rightarrow\infty,$
proved the quantitative lower bound (2.2), and shown that at least one component or both diverges under (1.2)-(2.3). This completes the proof of Theorem 2.3. ∎
4 Conclusion
In this work, we have extended the existence theory for entire large solutions of competitive semilinear elliptic systems beyond the classical power-law setting studied by Lair [3]. Our generalization applies to a broad class of nonlinearities that satisfy a Keller–Osserman-type growth condition, thereby encompassing both critical and supercritical regimes within a unified analytical framework.
The proof strategy preserves the essential structure of Lair’s original approach, while adapting each step to handle the greater generality of the nonlinear terms:
  • We employ a monotone iteration scheme to construct a global positive radial solution starting from admissible central values.
  • By summing the two equations, we reduce the system to a scalar differential inequality for $W=u+v$, which captures the combined growth of both components.
  • We apply a Keller–Osserman transform to $W$, exploiting the monotonicity and convexity properties of the transform to control the solution’s behaviour.
  • A two-step radial integration argument then forces blow-up of $W$ at infinity, ensuring that at least one component or both of the solution diverges.
This approach not only recovers the known results for pure power nonlinearities as a special case, but also provides a flexible method applicable to more general growth laws. In particular, the Keller–Osserman framework allows us to treat critical and supercritical growth in a single, coherent setting, highlighting the structural mechanisms that drive the existence of large solutions in competitive elliptic systems.