### v2

# Entire Large Solutions for Competitive Semilinear Elliptic Systems with General Nonlinearities Satisfying Keller–Osserman Conditions

Abstract
We generalize a theorem of Lair concerning the existence of positive entire large solutions to competitive semilinear elliptic systems. While Lair’s original result [3] was established for power-type nonlinearities, our work extends the theory to a broad class of general nonlinearities satisfying a Keller–Osserman-type growth condition. The proof follows the same conceptual framework monotone iteration to construct global positive solutions, reduction to a scalar inequality for the sum of the components, application of a Keller–Osserman transform, and a two-step radial integration argument but replaces the explicit power-law growth with a general monotone envelope function. This approach yields a unified and verifiable criterion for the existence of large solutions in terms of the Keller–Osserman integral, thereby encompassing both critical and supercritical growth regimes within a single analytical setting.
MSC 2020 Subject Classification: 35J47, 35J57, 35B40, 35B44.
Keywords and Phrases: Semilinear elliptic systems; entire large solutions; Keller–Osserman condition; monotone iteration; radial solutions; blow-up behaviour; competitive systems; nonlinear growth conditions.
1 Introduction
The search for positive entire large solutions to nonlinear elliptic systems is a classical theme in analysis, with deep connections to geometry, physics, and population models. In this work, we focus on the radial semilinear system
$\begin{cases}\Delta u=p(r)f(u,v),\\ \Delta v=q(r)g(u,v),\end{cases}\quad r=|x|,\quad x\in\mathbb{R}^{n},\ n\geq 3,$ (1.1)
We assume:
  • (H1)
    Weights: $p,q\geq 0$ are continuous, not both trivial, $\min(p,q)$ not compactly supported, with finite Green-type potentials
    $P(\infty)=\frac{1}{n-2}\int_{0}^{\infty}r\,p(r)\,dr<\infty,\quad Q(\infty)=\frac{1}{n-2}\int_{0}^{\infty}r\,q(r)\,dr<\infty,$
    where
    $P(r):=\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,ds\,dt,\quad Q(r):=\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,ds\,dt,$
    and
    $P(\infty)=\lim_{r\rightarrow\infty}P(r),Q(\infty)=\lim_{r\rightarrow\infty}Q(r).$
  • (H2)
    Monotonicity: $f,g$ are continuous on $0,\infty)^{2}$, $C^{1}$ on $(0,\infty)^{2}$, positive on $(0,\infty)^{2}$, vanish at $(0,0)$, and are nondecreasing in each variable.
  • (H3)
    Keller–Osserman envelope: There exists a continuous, nondecreasing $h\geq 0$, with $h(s)>0$ for $s>0$, such that for some small $\eta>0$,
    $f(s,t)+g(s,t)\geq h(s+t),\quad\forall s,t\geq\eta,$
    and the Keller–Osserman integral
    $H(\infty)=\lim_{s\rightarrow\infty}\int_{1}^{s}\frac{dt}{\sqrt{\int_{0}^{t}h(z)\,dz}}<\infty$
    is finite.
Our aim is to generalize the existence theory of Lair [3] "originally for power-type nonlinearities" to a broad class of nonlinearities governed by the Keller–Osserman growth condition. This unifies the treatment of critical and supercritical regimes, and recovers as special cases the results of Keller [2], Osserman [5], Cîrstea and Rădulescu [1], among others.
We call $(u,v)$ a positive entire solution if it is $C^{2}$, radial, and strictly positive on $0,\infty)$. The set of central values
[…]
encodes the initial data leading to such solutions. We show that $\mathcal{G}$ has a monotone structure, certain subsets are bounded and closed, and that on its boundary, all solutions are large at least one component diverges as $r\rightarrow\infty$.
Our method adapts Lair’s barrier and radial integration techniques [4, 3], replacing explicit power laws with a general monotone envelope $h$ satisfying the Keller–Osserman condition. This approach not only recovers known power-type results but also links naturally to complementary bounded-solution theories [8, 6, 7].
Structure of the paper. Section 2 states the main results. Section 3 contains the proofs, built on monotone iteration, Keller–Osserman estimates, and barrier arguments. Section 4 concludes with remarks and perspectives.
2 Main Results
We begin with the following property of the set $\mathcal{G}$.
Theorem 2.1.
Let $\alpha_{0}>0$, $\beta_{0}>0$. Then $\mathcal{G}$ is nonempty and has the property that if $(\alpha_{0},\beta_{0})\in\mathcal{G}$ and $0<\alpha<\alpha_{0}$, $0<\beta<\beta_{0}$, then $(\alpha,\beta)\in\mathcal{G}$.
For $\delta>0$ sufficiently small, define
$\mathcal{G}_{\delta}:=\left\{(\alpha,\beta)\in\mathcal{G}\ \middle|\ \min\{\alpha,\beta\}\geq\delta\right\}.$
We have the following result.
Theorem 2.2.
The set $\mathcal{G}_{\delta}$ is both closed and bounded in $(0,\infty)\times(0,\infty)$.
Next, we define the edge set
$\mathcal{E}(\mathcal{G}_{\delta}):=\left\{(\alpha,\beta)\in\partial\mathcal{G}_{\delta}\ \middle|\ \min\{\alpha,\beta\}>\delta\right\}.$
The geometric relationship between these sets is depicted in Figure 1, where $\mathcal{G}$ is shown in green, $\mathcal{G}_{\delta}$ in yellow, and the interior boundary arc $\mathcal{E}(\mathcal{G}_{\delta})$ in blue.
[…]
Figure 1: Existence region in the $(\alpha,\beta)$-plane: $G$ (green), $G_{\delta}$ (yellow), and $E(G_{\delta})$ (blue).
We can now state our main existence theorem for entire large solutions of (1.1).
Theorem 2.3.
There exists a positive radial solution $(u,v)$ of (1.1) with central value $(u(0),v(0))\in\mathcal{E}(\mathcal{G}_{\delta})$ such that at least one of $u$ or $v$ diverges, i.e.,
$\lim_{r\rightarrow\infty}u(r)=\infty\quad\text{or}\quad\lim_{r\rightarrow\infty}v(r)=\infty.$ (2.1)
Moreover, if $H(s):=\int_{s}^{\infty}\frac{dt}{f(t,t)+g\left(t,t\right)}$ and $H^{-1}$ denotes its inverse, then for all sufficiently large $r$,
$u(r)+v(r)\ \geq\ H^{-1}\!\left(\int_{r}^{\infty}t^{1-n}\int_{0}^{t}s^{n-1}\big(p(s)+q(s)\big)\,ds\,dt\right).$ (2.2)
If, in addition, we impose the following Keller-Osserman condition at infinity:
[…]
then $\lim_{r\rightarrow\infty}u(r)=\lim_{r\rightarrow\infty}v(r)=\infty$.
Remark 2.4 (Relation to Lair’s result).
When $f(u,v)=u^{a}v^{b}$ and $g(u,v)=u^{c}v^{d}$ with $\min\{a,d\}>1$, the envelope $h(s)=s^{\min\{a+b,c+d\}}$ satisfies the Keller–Osserman condition if and only if $\min\{a+b,c+d\}>1$. This recovers the main theorem of [3] as a special case.
Remark 2.5 (Relation to Zhang et al.’s result).
Both Theorem 2.3 and Theorem 1.2 of [8] study radial elliptic systems of the form (1.1). The two results are therefore compatible: they describe different behaviours (bounded vs. large) in disjoint regions of the central value space, even for the same coefficients.
In Theorem 1.2 of [8], the focus is on positive bounded entire radial solutions, under a strict hierarchy between the finite numbers $P(\infty)$, $Q(\infty)$ and $H(\infty)$.
Theorem 2.3 asserts that for $(u(0),v(0))\in\mathcal{E}(\mathcal{G}_{\delta})$, any entire solution is large, while Theorem 1.2 of [8] produces bounded entire solutions for $(u(0),v(0))$ in the interior of $\mathcal{G}$.
Remark 2.6.
The sets $\mathcal{G}$, $\mathcal{G}_{\delta}$, and $\mathcal{E}(\mathcal{G}_{\delta})$ can be thought of as nested regions: $\mathcal{G}$ contains all starting values that lead to positive solutions, $\mathcal{G}_{\delta}$ is a smaller core within it, and $\mathcal{E}(\mathcal{G}_{\delta})$ marks the area where the solutions grow without bound.
3 Proof of the Theorems
We recall that any radial $C^{2}$ solution $(u,v)$ of (1.1) satisfies the equivalent Volterra-type integral system
$u(r)$ $=\alpha_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,f\big(u(s),v(s)\big)\,ds\,dt,$ (3.1)
$v(r)$ $=\beta_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,g\big(u(s),v(s)\big)\,ds\,dt,$ (3.2)
for all $r\geq 0$, where $\alpha_{0}=u(0)$ and $\beta_{0}=v(0)$ are the central values. For $\alpha_{0},\beta_{0}\in(0,\infty)$, any solution of (3.1)–(3.2) is also a solution of (1.1), though it may not satisfy the large-solution condition (2.1).
3.1 Proof of Theorem 2.1
The proof of Theorem 2.1 is organized into six logically steps. We adapt the monotone iteration method of [3] to the present setting, where $f$ and $g$ satisfy the general hypotheses (H2)–(H3).
Step 1: Iterative scheme.
Proof.
Given $(\alpha_{0},\beta_{0})\in(0,\infty)\times(0,\infty)$, define
$u_{0}(r)\equiv\alpha_{0},\quad v_{0}(r)\equiv\beta_{0},$
and for $k\geq 1$ set
$u_{k}(r)$ $=\alpha_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,f\big(u_{k-1}(s),v_{k-1}(s)\big)\,ds\,dt,$
$v_{k}(r)$ $=\beta_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,g\big(u_{k-1}(s),v_{k-1}(s)\big)\,ds\,dt.$
Step 2: Monotonicity.
Proof.
By the positivity of $p,q$ and the monotonicity of $f,g$ in each variable (assumption (H2)), we have
$u_{k}(r)\geq u_{k-1}(r),\quad v_{k}(r)\geq v_{k-1}(r),\quad\forall r\geq 0,\ \forall k\geq 1.$
Thus, for each fixed $r$, both sequences $\{u_{k}(r)\}$ and $\{v_{k}(r)\}$ are nondecreasing. ∎
Step 3: Uniform upper bound via a scalar supersolution.
Proof.
Construction of $w$. The Keller–Osserman condition (H3) ensures the existence of a scalar function $w\in C^{2}([0,\infty))$ satisfying
$\Delta w=(p+q)\big(f(w,w)+g(w,w)\big),\quad w(0)>\alpha_{0}+\beta_{0},\quad\lim_{r\rightarrow\infty}w(r)=\infty.$ (3.3)
To construct $w$, consider the scalar radial ODE
$W^{\prime\prime}(r)+\frac{n-1}{r}W^{\prime}(r)=(p(r)+q(r))\,S(W(r)),$
where $S(s):=f(s,s)+g(s,s)$. By (H2)–(H3), $H$ is continuous, nondecreasing, positive for $s>0$, and satisfies the Keller–Osserman integral condition
$\int_{1}^{\infty}\frac{ds}{\sqrt{\int_{0}^{s}S(\tau)\,d\tau}}<\infty.$
Choosing $W(0)$ sufficiently large, standard ODE theory and the above integral condition guarantee that $W(r)\rightarrow\infty$ as $r\rightarrow\infty$. Setting $w:=W$ yields (3.3).
Significance of $w$. The function $w$ is a scalar supersolution for the system: it bounds the sum $u+v$ from above and grows unboundedly at infinity. This is crucial because:
  • It provides a uniform a priori bound on the iterates $u_{k}+v_{k}$, independent of $k$, on every compact interval.
  • It allows the use of comparison principles to control the nonlinear coupling in the system by a single scalar equation.
  • It ensures that the iteration cannot “blow up” prematurely on finite intervals.
By the comparison principle for radial integral equations, we have
$u_{k}(r)+v_{k}(r)\leq w(r),\quad\forall r\geq 0,\ \forall k\geq 0.$
Step 4: Convergence of the iteration.
Proof.
For each fixed $r\geq 0$, the sequences $\{u_{k}(r)\}$ and $\{v_{k}(r)\}$ are monotone increasing and bounded above by $w(r)$. Thus, the pointwise limits
$u(r):=\lim_{k\rightarrow\infty}u_{k}(r),\quad v(r):=\lim_{k\rightarrow\infty}v_{k}(r)$
exist for all $r\geq 0$.
Monotone Convergence Theorem (MCT). The MCT states: If $\{s_{k}\}$ is a sequence of nonnegative measurable functions such that $s_{k}(x)\uparrow s(x)$ pointwise, then
$\lim_{k\rightarrow\infty}\int s_{k}=\int\lim_{k\rightarrow\infty}s_{k}.$
In our setting, $s_{k}$ corresponds to the integrands in (3.1)–(3.2). Since $u_{k},v_{k}$ increase pointwise and $p,q,f,g$ are nonnegative, the integrands are nonnegative and increasing in $k$. Thus, MCT justifies passing the limit $k\rightarrow\infty$ inside the integrals, showing that $(u,v)$ satisfies (3.1)–(3.2). ∎
Step 5: Conclusion.
Proof.
The limit pair $(u,v)$ is a positive $C^{2}$ radial solution of (1.1) with $(u(0),v(0))=(\alpha_{0},\beta_{0})$. Thus $(\alpha_{0},\beta_{0})\in\mathcal{G}$, proving that $\mathcal{G}$ is nonempty. ∎
Step 6: Downward closure property.
Proof.
If $(\alpha_{0},\beta_{0})\in\mathcal{G}$ and $0<\alpha<\alpha_{0}$, $0<\beta<\beta_{0}$, then starting the iteration from $(\alpha,\beta)$ yields smaller initial data. By the same monotonicity and boundedness arguments, the iteration converges to a positive entire solution, hence $(\alpha,\beta)\in\mathcal{G}$. ∎
3.2 Proof of Theorem 2.2
We divide the proof into two logically independent parts: (i) boundedness of $\mathcal{G}_{\delta}$, and (ii) closedness of $\mathcal{G}_{\delta}$.
Part I: Boundedness of $\mathcal{G}_{\delta}$
Proof.
Step 0: Contradiction setup. Suppose, for contradiction, that $\mathcal{G}_{\delta}$ is unbounded in $(0,\infty)\times(0,\infty)$. From Theorem 2.1, we know that $\mathcal{G}$ is downward closed in each coordinate: if $(\alpha_{0},\beta_{0})\in\mathcal{G}$ and $0<\alpha<\alpha_{0}$, $0<\beta<\beta_{0}$, then $(\alpha,\beta)\in\mathcal{G}$.
In particular, if $(\alpha_{0},\beta_{0})\in\mathcal{G}_{\delta}$, then for any $(\alpha,\beta)$ with
$\delta\leq\alpha\leq\alpha_{0},\quad\delta\leq\beta\leq\beta_{0},$
we also have $(\alpha,\beta)\in\mathcal{G}_{\delta}$. This implies that if $\mathcal{G}_{\delta}$ were unbounded, it would necessarily contain an unbounded subset of the form
$\delta,\infty)\times\{\delta\}\quad\text{or}\quad\{\delta\}\times[\delta,\infty).$
Without loss of generality, assume
$\{\delta\}\times[\delta,\infty)\subseteq\mathcal{G}_{\delta}.$
Step 1: Choice of a scalar comparison function. Define
$\widetilde{p}(r):=\min\{p(r),\,q(r)\}.$
By assumption (H1), $\widetilde{p}$ is continuous, radial, and not compactly supported. Hence there exists $R>0$ such that $\widetilde{p}(R)>0$. From [4] (and the Keller–Osserman framework), there exists a large nonnegative radial solution $w$ of
$\Delta w=\widetilde{p}(r)\,h(w)\quad\text{in }[0,R),\quad w(0)=\delta,\quad\lim_{r\rightarrow R^{-}}w(r)=\infty.$
Here large means that $w$ blows up at the finite radius $R$.
Significance of $w$:
The function $w$ is a scalar supersolution for the coupled system: it satisfies the same type of equation as $u+v$ but with the minimal weight $\widetilde{p}$ and the scalar nonlinearity $h$. It will serve as a barrier function in the comparison argument: if $u+v$ starts above $w$ at $r=0$, the maximum principle forces $u+v$ to stay above $w$ until blow-up, which is impossible if $u+v$ is smooth on $[0,R)$.
Step 2: Construction of a candidate solution in $\mathcal{G}_{\delta}$. Since $\{\delta\}\times[\delta,\infty)\subseteq\mathcal{G}_{\delta}$, pick $\alpha_{0}>\delta$ and consider the initial data
$u(0)=\alpha_{0}>w(0)=\delta,\quad v(0)=\delta.$
By Theorem 2.1, there exists a positive entire solution $(u,v)$ of (3.1)–(3.2) with these initial values:
$u(r)$ $=\alpha_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,f(u(s),v(s))\,ds\,dt,$
$v(r)$ $=\delta+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,g(u(s),v(s))\,ds\,dt,$
for $0\leq r<\infty$.
Step 3: Differential inequality for $u+v$. Since $u(r)\geq\alpha_{0}>\delta$ and $v(r)\geq\delta$ for all $r\geq 0$, we have
$u(r)+v(r)\geq\alpha_{0}+\delta>2\delta.$
By assumption (H3), $f(s,t)+g(s,t)\geq h(s+t)$ for $s,t\geq\eta$ (and here $\eta\leq\delta$), so
$\Delta(u+v)=p(r)f(u,v)+q(r)g(u,v)\geq\widetilde{p}(r)\,h(u+v)\quad\text{in }[0,R).$
Step 4: Comparison with $w$ and contradiction. The function $w$ satisfies $\Delta w=\widetilde{p}(r)\,h(w)$ in $[0,R)$, $w(0)=\delta$, and blows up at $r=R$. By the strong maximum principle applied to u+vwu+v-w, if $u(0)+v(0)>w(0)$, then $u+v>w$ near $r=0$. However, since $w$ blows up at $R$ while $u+v$ is $C^{2}$ on $[0,R)$, the comparison principle forces $u+v\leq w$ in $[0,R)$ a contradiction with $u(0)+v(0)>w(0)$.
Therefore, $\mathcal{G}_{\delta}$ cannot be unbounded; it is bounded. ∎
Part II: Closedness of $\mathcal{G}_{\delta}$
Proof.
Step 1: Sequential closure argument. Let $\{(\alpha_{m},\beta_{m})\}\subset\mathcal{G}_{\delta}$ be a sequence converging to $(\alpha,\beta)\in\overline{\mathcal{G}_{\delta}}$. For each $m$, Theorem 2.1 guarantees the existence of a positive entire solution $(u_{m},v_{m})$ of (3.1)–(3.2) with
$u_{m}(0)=\alpha_{m},\quad v_{m}(0)=\beta_{m}.$
Step 2: Uniform bounds and equicontinuity. From the monotone iteration scheme in Theorem 2.1 and the Keller–Osserman supersolution $w$ from (3.3) with $w(0)>\alpha_{m}+\beta_{m}$, we deduce that the family $\{(u_{m},v_{m})\}$ is uniformly bounded on compact intervals and equicontinuous.
Step 3: Compactness and passage to the limit. By the Arzelá –Ascoli theorem, there exists a subsequence converging uniformly on compact sets to a limit $(u,v)$. Passing to the limit in (3.1)–(3.2) (justified by the dominated convergence theorem, since the integrands are continuous and uniformly bounded on compacts) shows that $(u,v)$ is a positive entire solution with
$u(0)=\alpha,\quad v(0)=\beta.$
Since $\min\{\alpha,\beta\}\geq\delta$ by construction, we have $(\alpha,\beta)\in\mathcal{G}_{\delta}$.
Conclusion: $\mathcal{G}_{\delta}$ is both bounded and closed. ∎
3.3 Proof of Theorem 2.3
For clarity, we structure the proof of Theorem 2.3 into seven distinct and logically self-contained steps.
Standing assumptions and notation.
Proof.
We work under hypotheses (H1)–(H3) and consider the radial system (1.1) in its equivalent integral (Green) form (3.1)–(3.2), with prescribed central values
$(u(0),v(0))=(\alpha,\beta)\in\mathcal{E}(\mathcal{G}_{\delta}).$
We introduce the Keller–Osserman transform
$H(s):=\int_{s}^{\infty}\frac{dt}{f(t,t)+g(t,t)},\quad s>0.$
By construction:
  • $H$ is strictly decreasing, with
    $H^{\prime}(s)=-\frac{1}{f(s,s)+g(s,s)}<0;$
  • $H^{\prime\prime}(s)\geq 0$ since $s\mapsto f(s,s)+g(s,s)$ is nondecreasing by (H2).
Step 1: Perturbation of initial data and blow-up for approximants
Proof.
From Theorem 2.1, $\mathcal{G}$ is nonempty and downward closed in each coordinate. From Theorem 2.2, $\mathcal{G}_{\delta}$ is closed and bounded for $\delta>0$ sufficiently small.
Let $(\alpha,\beta)\in\mathcal{E}(\mathcal{G}_{\delta})$. For each $k\geq 1$, consider the perturbed initial data
$\left(\alpha+\frac{1}{k},\ \beta+\frac{1}{k}\right).$
If this pair belonged to $\mathcal{G}$, then by downward closure and the definition of $\mathcal{G}_{\delta}$, $(\alpha,\beta)$ would lie in the interior of $\mathcal{G}_{\delta}$, contradicting $(\alpha,\beta)\in\mathcal{E}(\mathcal{G}_{\delta})$. Hence
$\left(\alpha+\frac{1}{k},\ \beta+\frac{1}{k}\right)\notin\mathcal{G}.$
Therefore, the solution $(u_{k},v_{k})$ of
$u_{k}(r)$ $=\alpha+\frac{1}{k}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)f(u_{k}(s),v_{k}(s))\,ds\,dt,$
$v_{k}(r)$ $=\beta+\frac{1}{k}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)g(u_{k}(s),v_{k}(s))\,ds\,dt,$
must blow up in finite radius: there exists $R_{k}>0$ such that
$\lim_{r\rightarrow R_{k}^{-}}u_{k}(r)=\infty\quad\text{or}\quad\lim_{r\rightarrow R_{k}^{-}}v_{k}(r)=\infty.$
In either case,
$\lim_{r\rightarrow R_{k}^{-}}\big(u_{k}(r)+v_{k}(r)\big)=\infty.$ (3.4)
Moreover, by monotonicity in $k$ of the blow-up radii (see [1]),
$R_{1}\leq R_{2}\leq\dots\leq R_{k}\leq\dots,$
and we set
$R:=\lim_{k\rightarrow\infty}R_{k}\in[R_{1},\infty].$
Step 2: Reduction to a scalar inequality
Proof.
Let
$W_{k}(r):=u_{k}(r)+v_{k}(r),\quad W(r):=u(r)+v(r),$
where $(u,v)$ is the limit solution corresponding to $(\alpha,\beta)$. Since $u_{k}\leq W_{k}$ and $v_{k}\leq W_{k}$, and $f,g$ are nondecreasing, we have
$\Delta W_{k}$ $=p(r)f(u_{k},v_{k})+q(r)g(u_{k},v_{k})$
$\leq(p(r)+q(r))\,\big(f(W_{k},W_{k})+g(W_{k},W_{k})\big),$ (3.5)
by (H2). ∎
Step 3: Applying the Keller–Osserman transform
Proof.
Applying $H$ to $W_{k}$ and using $H^{\prime\prime}\geq 0$, the chain rule gives
$\Delta H(W_{k})=H^{\prime}(W_{k})\,\Delta W_{k}+H^{\prime\prime}(W_{k})|\nabla W_{k}|^{2}\geq H^{\prime}(W_{k})\,\Delta W_{k}.$
Since $H^{\prime}(s)=-1/(f(s,s)+g(s,s))$, inequality (3.5) yields
$\Delta H(W_{k})\geq-(p(r)+q(r)).$
Step 4: First radial integration
Proof.
In radial coordinates, the above inequality becomes
$\big(r^{n-1}(H(W_{k}))^{\prime}(r)\big)^{\prime}\geq-r^{n-1}(p(r)+q(r)).$
Integrating from 0 to $r<R_{k}$ and using regularity at $r=0$ (which implies $\lim_{t\downarrow 0}t^{n-1}(H(W_{k}))^{\prime}(t)=0$), we obtain
$(H(W_{k}))^{\prime}\geq-r^{1-n}\int_{0}^{r}s^{n-1}(p(s)+q(s))\,ds.$
Step 5: Second radial integration and lower bound
Proof.
Integrating from $r$ to $R_{k}$ gives
$H(W_{k}(R_{k}))-H(W_{k}(r))\geq-\int_{r}^{R_{k}}t^{1-n}\left(\int_{0}^{t}s^{n-1}(p(s)+q(s))\,ds\right)dt.$
Letting $r\rightarrow R_{k}^{-}$, we have
$W_{k}(R_{k})\rightarrow\infty\text{ and hence }H(W_{k}(R_{k}))\rightarrow 0.$
Thus
$H(W_{k}(r))\leq\int_{r}^{R_{k}}t^{1-n}\left(\int_{0}^{t}s^{n-1}(p(s)+q(s))\,ds\right)dt.$
Since $H$ is strictly decreasing, we invert to obtain
$W_{k}(r)\geq H^{-1}\!\left(\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}(p(s)+q(s))\,ds\,dt\right).$
Passing to the limit $k\rightarrow\infty$ yields
$W(r)\geq H^{-1}\!\left(\int_{r}^{R}t^{1-n}\int_{0}^{t}s^{n-1}(p(s)+q(s))\,ds\,dt\right).$
If $R<\infty$, the right-hand side tends to $\infty$ as $r\rightarrow R^{-}$, contradicting the fact that $(u,v)$ is entire. Hence $R=\infty$, and we have established the quantitative lower bound (2.2). ∎
Step 6: Divergence of at least one component.
Proof.
From the previous steps (in particular, Step 1 of the proof of Theorem 2.2 and (3.4)), we already know that
$W(r):=u(r)+v(r)\;\longrightarrow\;+\infty\quad\text{as }r\rightarrow\infty.$
Suppose, for contradiction, that both $u$ and $v$ are bounded on $0,\infty)$. That is, there exist constants $M_{1},M_{2}>0$ such that
$u(r)\leq M_{1},\quad v(r)\leq M_{2}\quad\forall r\geq 0.$
Then
$W(r)\leq M_{1}+M_{2}\quad\forall r\geq 0,$
which contradicts $W(r)\rightarrow+\infty$. Therefore, at least one of the two components must diverge:
$\lim_{r\rightarrow\infty}u(r)=+\infty\quad\text{or}\quad\lim_{r\rightarrow\infty}v(r)=+\infty.$
This next figure shows
$u\left(r\right)\text{ and }v\left(r\right)\text{ vs. }r,$
with the “both bounded” scenario impossible, and the Keller–Osserman case forcing both curves upward.
[…]
Figure 2: $W(r)\rightarrow+\infty$ forces at least one component to diverge; under the Keller–Osserman condition both diverge. Solid lines: without KO; dashed lines: with KO.
Both components diverge under the supplementary Keller–Osserman condition. Assume now that the additional Keller–Osserman condition at infinity (2.3) holds. We claim that in this case $\lim_{r\rightarrow\infty}u(r)\;=\;\lim_{r\rightarrow\infty}v(r)\;=\;+\infty$. Indeed, if $u$ were bounded, then $v=W-u\rightarrow+\infty$. Thus it suffices to prove the implications
$v(r)\rightarrow+\infty\ \Longrightarrow\ u(r)\rightarrow+\infty,\qquad u(r)\rightarrow+\infty\ \Longrightarrow\ v(r)\rightarrow+\infty.$
Claim 1: If $v(r)\rightarrow+\infty$ as $r\rightarrow\infty$, then $u(r)\rightarrow+\infty$ as $r\rightarrow\infty$.
Proof of Claim 1. Assume by contradiction that $u$ is bounded: there exists $U>0$ such that
$0<u(r)\leq U\quad\forall r\geq 0.$
Since $v(r)\rightarrow+\infty$ while $u$ is bounded, there exists $R_{0}>0$ such that
$u(r)\leq U<v(r),\forall r\geq R_{0}.$
By the monotonicity of $f$ in each variable (assumption (H2)), we have
$f\big(u(r),v(r)\big)>f\big(u(r),u(r)\big)=:\ell_{f}\big(u(r)\big)\quad\forall r\geq R_{0},$
where $\ell_{f}(s):=f(s,s)$ is continuous, nondecreasing, satisfies $\ell_{f}(0)=0$, and, by hypothesis, the Keller–Osserman condition
$\int_{1}^{\infty}\frac{dt}{\sqrt{\int_{0}^{t}\ell_{f}(z)\,dz}}<\infty.$
Thus, for $r\geq R_{0}$,
$\Delta u(r)>p(r)\,\ell_{f}\big(u(r)\big).$ (3.6)
Barrier construction. Fix $R_{0}>0$ as above and consider the radial Cauchy problem
$y^{\prime\prime}(r)+\frac{n-1}{r}y^{\prime}(r)$ $=p(r)\,\ell_{f}\big(y(r)\big),\quad r\in[R_{0},R_{a}),$ (3.7)
$y(R_{0})$ $=a<u(R_{0}),\quad y^{\prime}(R_{0})=0,$
which has a maximal $C^{2}$ solution $y$ on $R_{0},R_{a})$ by standard ODE theory.
Define the Keller–Osserman primitive
$\mathcal{C}_{f}(s):=\int_{s_{\ast}}^{s}\frac{d\tau}{\ell_{f}(\tau)},\quad s>s_{\ast},$
where $s_{\ast}>0$ is chosen so that
$\ell_{f}(\tau)>0\text{ on }[s_{\ast},\infty).$
Then $\mathcal{C}_{f}$ is strictly increasing on $(s_{\ast},\infty)$, with
$\mathcal{C}_{f}^{\prime}(s)=\frac{1}{\ell_{f}(s)},\quad\mathcal{C}_{f}^{\prime\prime}(s)=-\frac{\ell_{f}^{\prime}(s)}{\ell_{f}(s)^{2}}\leq 0,$
and, by the Keller–Osserman condition, $\mathcal{C}_{f}(\infty)<\infty$.
Set $W(r):=\mathcal{C}_{f}\big(y(r)\big)$. Differentiating and using (3.7) yields
$W^{\prime}(r)=\frac{y^{\prime}(r)}{\ell_{f}(y(r))},\quad\big(r^{n-1}W^{\prime}(r)\big)^{\prime}=\frac{r^{n-1}}{\ell_{f}(y)}\left(y^{\prime\prime}+\frac{n-1}{r}y^{\prime}\right)-\frac{r^{n-1}\ell_{f}^{\prime}(y)}{\ell_{f}(y)^{2}}\,(y^{\prime})^{2}\leq r^{n-1}p(r).$
Integrating from $R_{0}$ to $r$ and using $W^{\prime}(R_{0})=y^{\prime}\left(R_{0}\right)/\mathcal{C}_{f}(a)=0$ gives
$r^{n-1}W^{\prime}(r)\leq\int_{R_{0}}^{r}s^{n-1}p(s)\,ds.$
Dividing by $r^{n-1}$ and integrating again from $R_{0}$ to $r$ yields the Keller–Osserman estimate
$\mathcal{C}_{f}\big(y(r)\big)\leq\mathcal{C}_{f}(a)+Q_{R_{0}}(r),\quad Q_{R_{0}}(r):=\int_{R_{0}}^{r}t^{1-n}\left(\int_{R_{0}}^{t}s^{n-1}p(s)\,ds\right)dt.$ (3.8)
By direct computation (Fubini’s theorem and $n\geq 3$),
$0\leq Q_{R_{0}}(\infty)=\frac{1}{n-2}\int_{R_{0}}^{\infty}\left(1-\left(\frac{R_{0}}{t}\right)^{n-2}\right)t\,p(t)\,dt\leq\frac{1}{n-2}\int_{R_{0}}^{\infty}t\,p(t)\,dt,$
so $Q_{R_{0}}(\infty)<\infty$ and $Q_{R_{0}}(\infty)\downarrow 0$ as $R_{0}\uparrow\infty$.
Blow-up of the barrier. Since $\mathcal{C}_{f}(\infty)<\infty$, we can first choose $R_{0}$ large so that $Q_{R_{0}}(\infty)$ is as small as desired, then choose $a$ close enough to $u(R_{0})$ but still smaller than $u(R_{0})$, so that $\mathcal{C}_{f}(a)\;+\;Q_{R_{0}}(\infty)\;>\;\mathcal{C}_{f}(\infty)$. Because $Q_{R_{0}}$ is continuous and strictly increasing, there exists a finite $R_{a}\in(R_{0},\infty)$ with
$\mathcal{C}_{f}(a)\;+\;Q_{R_{0}}(R_{a})\;=\;\mathcal{C}_{f}(\infty).$ (3.9)
By (3.8), this forces $\mathcal{C}_{f}\bigl(y(r)\bigr)\nearrow\mathcal{C}_{f}(\infty)$ as $r\uparrow R_{a}$, hence $\lim_{r\uparrow R_{a}}y(r)\;=\;+\infty$, i.e., $y$ is a (finite-radius) blow-up supersolution for (3.6) on $R_{0},R_{a})$.
Maximum principle comparison and contradiction. Fix any $R\in(R_{0},R_{a})$. Since $u$ is assumed bounded on $R_{0},\infty)$ while $y(r)\rightarrow\infty$ as $r\uparrow R_{a}$, we may choose $R$ sufficiently close to $R_{a}$ so that
$y(R)>\sup_{r\geq R_{0}}u(r).$
In particular,
$u(R_{0})-y(R_{0})=u(R_{0})-a>0,\qquad u(R)-y(R)<0.$
. enDefine w:=uyw:=u-y on the annulus
$A_{R_{0},R}:=\bigl\{x\in\mathbb{R}^{n}:\;R_{0}<r=|x|<R\bigr\}\,\}.$
Using (3.6) and (3.7), together with the monotonicity of $l_{f}$, we obtain on the set where $u\geq y$:
$\Delta w=\Delta u-\Delta y>p(r)\,l_{f}(u)-p(r)\,l_{f}(y)\geq 0.$
Let
$r_{\ast}:=\inf\{\,r\in(R_{0},R]:w(r)\leq 0\,\}.$
By the boundary values above, $r_{\ast}\in(R_{0},R)$ and
$w(r)>0\quad\text{for }r\in[R_{0},r_{\ast}),\qquad w(r_{\ast})=0.$
Thus $w$ is nonnegative on $R_{0},r_{\ast}]$ and strictly positive inside, with $w$ attaining its minimum 0 at the boundary point $r_{\ast}$ of the domain $(R_{0},r_{\ast})$. Moreover, on $(R_{0},r_{\ast})$ we have $u\geq y$ and hence $\Delta w>0$.
The situation is illustrated in Figure, where the blue curve denotes $u$, the red curve the barrier $y$.
[…]
Figure 3: Schematic of the first contact $r_{\ast}$ where $u$ (blue) and $y$ (red) intersect: solid for $r<r_{\ast}$, dashed for $r>r_{\ast}$. On $(R_{0},r_{\ast})$ we have w=uy>0w=u-y>0 and $\Delta w>0$, so the Hopf boundary lemma yields $w^{\prime}(r_{\ast})>0$, contradicting the first-contact requirement $w^{\prime}(r_{\ast})\leq 0$.
By the (radial) weak maximum principle applied to $w$ on $(R_{0},r_{\ast})$, $w$ cannot achieve a strict interior minimum; by the Hopf boundary point lemma at $r_{\ast}$, the outward normal derivative must satisfy
$w^{\prime}(r_{\ast})>0\text{ (outward normal points to ward increasing }r\text{)}.$
However, $r_{\ast}$ is the first contact point where $w$ crosses from positive to nonpositive values, so necessarily $w^{\prime}(r_{\ast})\leq 0$. This contradiction shows that such $r_{\ast}$ cannot exist. The next Figure shows $u$ vs. $y$ and why the “forbidden crossing” at $r_{\ast}$ can’t happen.
[…]
Figure 4: $u=y$ separates the regions $u>y$ and $u<y$. The trajectory starts in $u>y$ at $R_{0}$ and touches the diagonal for the first time at $r_{\ast}$. The maximum principle forbids crossing into $u<y$ here.
Hence w(r)=u(r)y(r)>0w(r)=u(r)-y(r)>0 on $R_{0},R)$. Letting $R\uparrow R_{a}$ yields $u(r)\geq y(r)$ for all $r\in[R_{0},R_{a})$. Since $y(r)\rightarrow\infty$ as $r\uparrow R_{a}$, we must have $u(r)\rightarrow\infty$ as $r\uparrow R_{a}$, which contradicts the assumed boundedness of $u$ (and contradicts the prescribed initial value comparison $u(R_{0})>a$ together with the smoothness of $u$ on $R_{0},\infty)$). Therefore, the assumption that $u$ is bounded while $v\rightarrow\infty$ is impossible.
Claim 2: If $u(r)\rightarrow+\infty$ as $r\rightarrow\infty$, then $v(r)\rightarrow+\infty$ as $r\rightarrow\infty$.
Proof of Claim 2. The argument is symmetric, using $\ell_{g}$, $q$, and a barrier $w$ with $w(R_{1})<v(R_{1})$ to obtain the same type of contradiction. ∎∎
Step 7: Conclusion.
Proof.
We have now completed the construction of a positive, radial, entire solution
$(u,v)\in C^{2}([0,\infty))\times C^{2}([0,\infty))$
to the system (1.1) under the standing assumptions (H1)–(H3). From the previous steps, we have established that the sum
$W(r):=u(r)+v(r)$
satisfies
$W(r)\longrightarrow+\infty\quad asr\rightarrow\infty,$
and moreover obeys the quantitative lower bound (2.2) derived via the Keller–Osserman transform.
We have also shown that:
  • Without any additional conditions added to (H1)-(H3), at least one of the two components $u$ or $v$ must diverge to $\infty$ as $r\rightarrow\infty$.
  • Under the supplementary Keller–Osserman condition at infinity (2.3), both components necessarily diverge:
    $\lim_{r\rightarrow\infty}u(r)=\lim_{r\rightarrow\infty}v(r)=+\infty.$
Thus, the constructed pair $(u,v)$ is a positive, radial, entire solution of (1.1) with the stated asymptotic properties, completing the proof of Theorem 2.3. ∎
4 Conclusion
In this work, we have extended the existence theory for entire large solutions of competitive semilinear elliptic systems beyond the classical power-law setting studied by Lair [3]. Our generalization applies to a broad class of nonlinearities that satisfy a Keller–Osserman-type growth condition, thereby encompassing both critical and supercritical regimes within a unified analytical framework.
The proof strategy preserves the essential structure of Lair’s original approach, while adapting each step to handle the greater generality of the nonlinear terms:
  • We employ a monotone iteration scheme to construct a global positive radial solution starting from admissible central values.
  • By summing the two equations, we reduce the system to a scalar differential inequality for $W=u+v$, which captures the combined growth of both components.
  • We apply a Keller–Osserman transform to $W$, exploiting the monotonicity and convexity properties of the transform to control the solution’s behaviour.
  • A two-step radial integration argument then forces blow-up of $W$ at infinity, ensuring that at least one component or both of the solution diverges.
This approach not only recovers the known results for pure power nonlinearities as a special case, but also provides a flexible method applicable to more general growth laws. In particular, the Keller–Osserman framework allows us to treat critical and supercritical growth in a single, coherent setting, highlighting the structural mechanisms that drive the existence of large solutions in competitive elliptic systems.