### v3

# Entire Large Solutions for Competitive Semilinear Elliptic Systems with General Nonlinearities Satisfying Keller–Osserman Conditions

Abstract
We establish the existence of positive entire large solutions for competitive semilinear elliptic systems with general nonlinearities, significantly extending recent work of Lair [3] from power-type to Keller–Osserman-type growth conditions. For the radial system $\Delta u=p(r)f(u,v)$, $\Delta v=q(r)g(u,v)$ in $\mathbb{R}^{n}$ ($n\geq 3$), we prove that under appropriate integrability conditions on the weights $p,q$ and a Keller–Osserman condition on the envelope function $h$ satisfying $f(s,t)+g(s,t)\geq h(s+t)$, there exists a positive radial solution $(u,v)$ with at least one component diverging to infinity. Our proof employs monotone iteration, scalar reduction via comparison principles, the Keller–Osserman transform, and barrier arguments. The main result provides explicit quantitative lower bounds on the blow-up rate and establishes conditions under which both components diverge simultaneously. This unified framework recovers classical power-law results as special cases while encompassing critical and supercritical growth regimes, and complements the bounded-solution theory in the literature.
MSC 2020 Subject Classification: 35J47, 35J57, 35B40, 35B44.
Keywords and Phrases: Semilinear elliptic systems; entire large solutions; Keller–Osserman condition; monotone iteration; radial solutions; blow-up behaviour; competitive systems; nonlinear growth conditions.
1 Introduction
The study of positive entire large solutions to nonlinear elliptic systems is a fundamental area in the theory of partial differential equations, with profound connections to differential geometry, mathematical physics, reaction-diffusion systems, and population dynamics. Solutions that blow up at infinity arise naturally in various contexts: prescribed scalar curvature problems [2, 5], conformal geometry, steady-state models of chemical reactions, and competitive or cooperative biological systems.
The existence and qualitative behavior of such large solutions have been investigated extensively since the pioneering works of Keller [2] and Osserman [5] in the 1950s, who established the now-classical Keller–Osserman condition for scalar equations. This condition characterizes when a single semilinear equation $\Delta u=f(u)$ admits solutions that diverge to infinity at spatial infinity. Extensions to systems began with the work of Cîrstea and Rădulescu [1], and have seen significant developments in recent years, including the work of Zhang et al. [8] on bounded solutions, Peterson and Wood [6] on non-monotone systems, and more recently, Lair [3] for competitive systems with power-type nonlinearities.
In this work, we focus on the radial semilinear system
$\begin{cases}\Delta u=p(r)f(u,v),\\ \Delta v=q(r)g(u,v),\end{cases}\quad r=|x|,\quad x\in\mathbb{R}^{n},\ n\geq 3,$ (1.1)
in the entire space $\mathbb{R}^{n}$, where the weight functions $p,q$ are radially symmetric and satisfy appropriate integrability conditions, and the nonlinearities $f,g$ satisfy general Keller–Osserman-type growth conditions.
We assume:
  • (H1)
    Weights: $p,q\geq 0$ are continuous, not both trivial, $\min(p,q)$ not compactly supported, with finite Green-type potentials
    $P(\infty)=\frac{1}{n-2}\int_{0}^{\infty}r\,p(r)\,dr<\infty,\quad Q(\infty)=\frac{1}{n-2}\int_{0}^{\infty}r\,q(r)\,dr<\infty,$
    where
    $P(r):=\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,ds\,dt,\quad Q(r):=\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,ds\,dt,$
    and
    $P(\infty)=\lim_{r\rightarrow\infty}P(r),Q(\infty)=\lim_{r\rightarrow\infty}Q(r).$
  • (H2)
    Monotonicity: $f,g$ are continuous on $0,\infty)^{2}$, $C^{1}$ on $(0,\infty)^{2}$, positive on $(0,\infty)^{2}$, vanish at $(0,0)$, and are nondecreasing in each variable.
  • (H3)
    Keller–Osserman envelope: There exists a continuous, nondecreasing $h\geq 0$, with $h(s)>0$ for $s>0$, such that for some small $\eta>0$,
    $f(s,t)+g(s,t)\geq h(s+t),\quad\forall s,t\geq\eta,$
    and the Keller–Osserman integral
    $H(\infty)=\lim_{s\rightarrow\infty}\int_{1}^{s}\frac{dt}{\sqrt{\int_{0}^{t}h(z)\,dz}}<\infty$
    is finite.
Main contributions. Our principal aim is to extend the existence theory of Lair [3], originally established for power-type nonlinearities, to a broad class of nonlinearities satisfying Keller–Osserman-type growth conditions. This extension is both natural and necessary:
  • (i)
    Generality: Many physical and geometric problems lead to nonlinearities that do not have pure power-law form. Our framework accommodates these more general growth conditions while preserving the essential analytical structure.
  • (ii)
    Unification: By working with a general envelope function $h$ satisfying the Keller–Osserman integral condition, we provide a unified treatment of critical and supercritical growth regimes, recovering as special cases not only Lair’s results but also classical results of Keller [2], Osserman [5], and Cîrstea and Rădulescu [1].
  • (iii)
    Completeness: Our results complement the bounded-solution theories developed by Zhang et al. [8], Peterson and Wood [6], and Santos et al. [7], providing a complete picture of the existence landscape for system (1.1).
Analytical framework. We call $(u,v)$ a positive entire solution if it is $C^{2}$, radial, and strictly positive on $0,\infty)$. A solution is called large if at least one component diverges to infinity as $r\to\infty$. The set of central values
[…]
encodes the initial data leading to such solutions. A key insight of our work is that $\mathcal{G}$ has a rich geometric structure: it is downward-closed (in the sense that smaller initial values also lead to entire solutions), certain regularized subsets $\mathcal{G}_{\delta}$ are compact, and solutions corresponding to boundary points of $\mathcal{G}_{\delta}$ are necessarily large—at least one component diverges as $r\rightarrow\infty$.
Methodology. Our approach adapts and extends Lair’s barrier and radial integration techniques [4, 3] to the setting of general nonlinearities. The key steps are:
  • Monotone iteration: We construct solutions via a monotone iterative scheme, with upper bounds provided by scalar supersolutions.
  • Scalar reduction: By exploiting the monotonicity of $f$ and $g$, we reduce the coupled system to a scalar differential inequality for $W=u+v$.
  • Keller–Osserman transform: We apply a nonlinear transformation $H$ to $W$, converting the nonlinear problem into a linear inequality amenable to explicit integration.
  • Blow-up dichotomy: Using barrier arguments and the Hopf lemma, we establish that solutions with boundary initial data must blow up.
This approach not only recovers known power-type results but also provides explicit quantitative bounds on the blow-up rate.
Organization. Section 2 states the main results. Section 3 contains the complete proofs, built on monotone iteration, Keller–Osserman estimates, and barrier arguments. Section 4 concludes with remarks and perspectives for future research.
2 Main Results
We begin with the following property of the set $\mathcal{G}$.
Theorem 2.1.
Let $\alpha_{0}>0$, $\beta_{0}>0$. Then $\mathcal{G}$ is nonempty and has the property that if $(\alpha_{0},\beta_{0})\in\mathcal{G}$ and $0<\alpha<\alpha_{0}$, $0<\beta<\beta_{0}$, then $(\alpha,\beta)\in\mathcal{G}$.
For $\delta>0$ sufficiently small, define
$\mathcal{G}_{\delta}:=\left\{(\alpha,\beta)\in\mathcal{G}\ \middle|\ \min\{\alpha,\beta\}\geq\delta\right\}.$
We have the following result.
Theorem 2.2.
The set $\mathcal{G}_{\delta}$ is both closed and bounded in $(0,\infty)\times(0,\infty)$.
Next, we define the edge set
$\mathcal{E}(\mathcal{G}_{\delta}):=\left\{(\alpha,\beta)\in\partial\mathcal{G}_{\delta}\ \middle|\ \min\{\alpha,\beta\}>\delta\right\}.$
The geometric relationship between these sets is depicted in Figure 1, where $\mathcal{G}$ is shown in green, $\mathcal{G}_{\delta}$ in yellow, and the interior boundary arc $\mathcal{E}(\mathcal{G}_{\delta})$ in blue.
$\alpha$$\beta$$\alpha=\delta$$\beta=\delta$$\mathcal{G}$$\mathcal{G}_{\delta}$$\mathcal{E}(\mathcal{G}_{\delta})$
Figure 1: Existence region in the $(\alpha,\beta)$-plane: $\mathcal{G}$ (green), $\mathcal{G}_{\delta}$ (yellow), and $\mathcal{E}(\mathcal{G}_{\delta})$ (blue).
We can now state our main existence theorem for entire large solutions of (1.1).
Theorem 2.3.
There exists a positive radial solution $(u,v)$ of (1.1) with central value $(u(0),v(0))\in\mathcal{E}(\mathcal{G}_{\delta})$ such that at least one of $u$ or $v$ diverges, i.e.,
$\lim_{r\rightarrow\infty}u(r)=\infty\quad\text{or}\quad\lim_{r\rightarrow\infty}v(r)=\infty.$ (2.1)
Moreover, if $H(s):=\int_{s}^{\infty}\frac{dt}{f(t,t)+g\left(t,t\right)}$ and $H^{-1}$ denotes its inverse, then for all sufficiently large $r$,
$u(r)+v(r)\ \geq\ H^{-1}\!\left(\int_{r}^{\infty}t^{1-n}\int_{0}^{t}s^{n-1}\big(p(s)+q(s)\big)\,ds\,dt\right).$ (2.2)
If, in addition, we impose the following Keller-Osserman condition at infinity:
[…]
then $\lim_{r\rightarrow\infty}u(r)=\lim_{r\rightarrow\infty}v(r)=\infty$.
Remark 2.4 (Relation to Lair’s result).
When $f(u,v)=u^{a}v^{b}$ and $g(u,v)=u^{c}v^{d}$ with $\min\{a,d\}>1$, the envelope $h(s)=s^{\min\{a+b,c+d\}}$ satisfies the Keller–Osserman condition if and only if $\min\{a+b,c+d\}>1$. This recovers the main theorem of [3] as a special case.
Remark 2.5 (Relation to Zhang et al.’s result).
Both Theorem 2.3 and Theorem 1.2 of [8] study radial elliptic systems of the form (1.1). The two results are therefore compatible: they describe different behaviours (bounded vs. large) in disjoint regions of the central value space, even for the same coefficients.
In Theorem 1.2 of [8], the focus is on positive bounded entire radial solutions, under a strict hierarchy between the finite numbers $P(\infty)$, $Q(\infty)$ and $H(\infty)$.
Theorem 2.3 asserts that for $(u(0),v(0))\in\mathcal{E}(\mathcal{G}_{\delta})$, any entire solution is large, while Theorem 1.2 of [8] produces bounded entire solutions for $(u(0),v(0))$ in the interior of $\mathcal{G}$.
Remark 2.6.
The sets $\mathcal{G}$, $\mathcal{G}_{\delta}$, and $\mathcal{E}(\mathcal{G}_{\delta})$ can be thought of as nested regions: $\mathcal{G}$ contains all starting values that lead to positive solutions, $\mathcal{G}_{\delta}$ is a smaller core within it, and $\mathcal{E}(\mathcal{G}_{\delta})$ marks the area where the solutions grow without bound.
3 Proof of the Theorems
We recall that any radial $C^{2}$ solution $(u,v)$ of (1.1) satisfies the equivalent Volterra-type integral system
$u(r)$ $=\alpha_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,f\big(u(s),v(s)\big)\,ds\,dt,$ (3.1)
$v(r)$ $=\beta_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,g\big(u(s),v(s)\big)\,ds\,dt,$ (3.2)
for all $r\geq 0$, where $\alpha_{0}=u(0)$ and $\beta_{0}=v(0)$ are the central values. For $\alpha_{0},\beta_{0}\in(0,\infty)$, any solution of (3.1)–(3.2) is also a solution of (1.1), though it may not satisfy the large-solution condition (2.1).
3.1 Proof of Theorem 2.1
Proof.
The proof is organized into six logical steps. We adapt the monotone iteration method of [3] to the present setting, where $f$ and $g$ satisfy the general hypotheses (H2)–(H3).
Step 1: Iterative scheme. Given $(\alpha_{0},\beta_{0})\in(0,\infty)\times(0,\infty)$, define
$u_{0}(r)\equiv\alpha_{0},\quad v_{0}(r)\equiv\beta_{0},$
and for $k\geq 1$ set
$u_{k}(r)$ $=\alpha_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,f\big(u_{k-1}(s),v_{k-1}(s)\big)\,ds\,dt,$
$v_{k}(r)$ $=\beta_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,g\big(u_{k-1}(s),v_{k-1}(s)\big)\,ds\,dt.$
Step 2: Monotonicity. By the positivity of $p,q$ and the monotonicity of $f,g$ in each variable (assumption (H2)), we have
$u_{k}(r)\geq u_{k-1}(r),\quad v_{k}(r)\geq v_{k-1}(r),\quad\forall r\geq 0,\ \forall k\geq 1.$
Thus, for each fixed $r$, both sequences $\{u_{k}(r)\}$ and $\{v_{k}(r)\}$ are nondecreasing.
Step 3: Uniform upper bound via a scalar supersolution. Construction of $w$. The Keller–Osserman condition (H3) ensures the existence of a scalar function $w\in C^{2}([0,\infty))$ satisfying
$\Delta w=(p+q)\big(f(w,w)+g(w,w)\big),\quad w(0)>\alpha_{0}+\beta_{0},\quad\lim_{r\rightarrow\infty}w(r)=\infty.$ (3.3)
To construct $w$ rigorously, consider the scalar radial ODE
$W^{\prime\prime}(r)+\frac{n-1}{r}W^{\prime}(r)=(p(r)+q(r))\,S(W(r)),$ (3.4)
where $S(s):=f(s,s)+g(s,s)$. By (H2)–(H3), $S$ is continuous, nondecreasing, positive for $s>0$, vanishes at $s=0$, and satisfies the Keller–Osserman integral condition
$\int_{1}^{\infty}\frac{ds}{\sqrt{\int_{0}^{s}S(\tau)\,d\tau}}<\infty.$ (3.5)
By classical existence theory for radial ODEs (see [2, 5]), for any initial value $W(0)=W_{0}>0$, there exists a maximal solution $W$ of (3.4) on an interval $0,R_{W_{0}})$ with $R_{W_{0}}\in(0,\infty]$. The Keller–Osserman condition (3.5) guarantees that if $W_{0}$ is chosen sufficiently large, then $R_{W_{0}}=\infty$ and $W(r)\to\infty$ as $r\to\infty$ (see [1], Theorem 2.1). Specifically, choosing $W_{0}>\alpha_{0}+\beta_{0}$, we obtain a global solution $W$ satisfying $W(0)=W_{0}>\alpha_{0}+\beta_{0}$ and $\lim_{r\to\infty}W(r)=\infty$. Setting $w:=W$ yields (3.3).
Significance of $w$. The function $w$ is a scalar supersolution for the system: it bounds the sum $u+v$ from above and grows unboundedly at infinity. This is crucial because:
  • It provides a uniform a priori bound on the iterates $u_{k}+v_{k}$, independent of $k$, on every compact interval.
  • It allows the use of comparison principles to control the nonlinear coupling in the system by a single scalar equation.
  • It ensures that the iteration cannot blow up prematurely on finite intervals.
By the comparison principle for radial integral equations (which follows from the maximum principle for the Laplacian), since $u_{0}(r)+v_{0}(r)=\alpha_{0}+\beta_{0}<w(0)$ and
$\Delta(u_{k}+v_{k})=p(r)f(u_{k},v_{k})+q(r)g(u_{k},v_{k})\leq(p(r)+q(r))\big(f(u_{k}+v_{k},u_{k}+v_{k})+g(u_{k}+v_{k},u_{k}+v_{k})\big),$
by monotonicity of $f,g$, we deduce by induction that
$u_{k}(r)+v_{k}(r)\leq w(r),\quad\forall r\geq 0,\ \forall k\geq 0.$
Step 4: Convergence of the iteration. For each fixed $r\geq 0$, the sequences $\{u_{k}(r)\}$ and $\{v_{k}(r)\}$ are monotone increasing (by Step 2) and bounded above by $w(r)$ (by Step 3). Thus, the pointwise limits
$u(r):=\lim_{k\rightarrow\infty}u_{k}(r),\quad v(r):=\lim_{k\rightarrow\infty}v_{k}(r)$
exist for all $r\geq 0$ by monotone convergence of real sequences.
Justification via the Monotone Convergence Theorem (MCT). To verify that $(u,v)$ satisfies the integral system (3.1)–(3.2), we apply Lebesgue’s MCT. The MCT states: If $\{\phi_{k}\}$ is a sequence of nonnegative measurable functions such that $\phi_{k}(x)\uparrow\phi(x)$ pointwise almost everywhere, then
$\lim_{k\rightarrow\infty}\int\phi_{k}\,d\mu=\int\lim_{k\rightarrow\infty}\phi_{k}\,d\mu=\int\phi\,d\mu.$
In our setting, fix $r>0$ and consider the innermost integral in (3.1):
$I_{k}(t):=\int_{0}^{t}s^{n-1}p(s)f(u_{k-1}(s),v_{k-1}(s))\,ds.$
Since $u_{k-1}(s)\uparrow u(s)$ and $v_{k-1}(s)\uparrow v(s)$ pointwise, and $f$ is nondecreasing and continuous, we have
$p(s)f(u_{k-1}(s),v_{k-1}(s))\uparrow p(s)f(u(s),v(s))\quad\text{pointwise in }s.$
The integrands are nonnegative. By MCT,
$\lim_{k\to\infty}I_{k}(t)=\int_{0}^{t}s^{n-1}p(s)f(u(s),v(s))\,ds.$
Moreover, $I_{k}(t)\uparrow$ and is bounded by
$I_{k}(t)\leq\int_{0}^{t}s^{n-1}(p(s)+q(s))S(w(s))\,ds=:M(t)<\infty$
for each $t<\infty$, by construction of $w$. Applying MCT again to the outer integral in (3.1), we obtain
$\lim_{k\to\infty}u_{k}(r)=\alpha_{0}+\int_{0}^{r}t^{1-n}\left(\lim_{k\to\infty}I_{k}(t)\right)dt=\alpha_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)f(u(s),v(s))\,ds\,dt.$
Thus $u=\lim_{k\to\infty}u_{k}$ satisfies (3.1). The same argument applies to $v$. By standard elliptic regularity, $(u,v)\in C^{2}([0,\infty))\times C^{2}([0,\infty))$.
Step 5: Conclusion. The limit pair $(u,v)$ is a positive $C^{2}$ radial solution of (1.1) with $(u(0),v(0))=(\alpha_{0},\beta_{0})$. Thus $(\alpha_{0},\beta_{0})\in\mathcal{G}$, proving that $\mathcal{G}$ is nonempty.
Step 6: Downward closure property. If $(\alpha_{0},\beta_{0})\in\mathcal{G}$ and $0<\alpha<\alpha_{0}$, $0<\beta<\beta_{0}$, then starting the iteration from $(\alpha,\beta)$ yields smaller initial data. By the same monotonicity and boundedness arguments (Steps 1–4), the iteration converges to a positive entire solution with initial values $(\alpha,\beta)$, hence $(\alpha,\beta)\in\mathcal{G}$. This completes the proof of Theorem 2.1. ∎
3.2 Proof of Theorem 2.2
Proof.
We divide the proof into two logically independent parts: (i) boundedness of $\mathcal{G}_{\delta}$, and (ii) closedness of $\mathcal{G}_{\delta}$.
Part I: Boundedness of $\mathcal{G}_{\delta}$.
Step 0: Contradiction setup. Suppose, for contradiction, that $\mathcal{G}_{\delta}$ is unbounded in $(0,\infty)\times(0,\infty)$. From Theorem 2.1, we know that $\mathcal{G}$ is downward closed in each coordinate: if $(\alpha_{0},\beta_{0})\in\mathcal{G}$ and $0<\alpha<\alpha_{0}$, $0<\beta<\beta_{0}$, then $(\alpha,\beta)\in\mathcal{G}$.
In particular, if $(\alpha_{0},\beta_{0})\in\mathcal{G}_{\delta}$, then for any $(\alpha,\beta)$ with
$\delta\leq\alpha\leq\alpha_{0},\quad\delta\leq\beta\leq\beta_{0},$
we also have $(\alpha,\beta)\in\mathcal{G}_{\delta}$. This implies that if $\mathcal{G}_{\delta}$ were unbounded, it would necessarily contain an unbounded subset of the form
$\delta,\infty)\times\{\delta\}\quad\text{or}\quad\{\delta\}\times[\delta,\infty).$
Without loss of generality, assume
$\{\delta\}\times[\delta,\infty)\subseteq\mathcal{G}_{\delta}.$
Step 1: Choice of a scalar comparison function. Define
$\widetilde{p}(r):=\min\{p(r),\,q(r)\}.$
By assumption (H1), $\widetilde{p}$ is continuous, radial, and not compactly supported. Hence there exists $R>0$ such that $\widetilde{p}(R)>0$. From [4] (and the Keller–Osserman framework), there exists a large nonnegative radial solution $w$ of
$\Delta w=\widetilde{p}(r)\,h(w)\quad\text{in }[0,R),\quad w(0)=\delta,\quad\lim_{r\rightarrow R^{-}}w(r)=\infty.$
Here large means that $w$ blows up at the finite radius $R$.
Significance of $w$:
The function $w$ is a scalar supersolution for the coupled system: it satisfies the same type of equation as $u+v$ but with the minimal weight $\widetilde{p}$ and the scalar nonlinearity $h$. It will serve as a barrier function in the comparison argument: if $u+v$ starts above $w$ at $r=0$, the maximum principle forces $u+v$ to stay above $w$ until blow-up, which is impossible if $u+v$ is smooth on $[0,R)$.
Step 2: Construction of a candidate solution in $\mathcal{G}_{\delta}$. Since $\{\delta\}\times[\delta,\infty)\subseteq\mathcal{G}_{\delta}$, pick $\alpha_{0}>\delta$ and consider the initial data
$u(0)=\alpha_{0}>w(0)=\delta,\quad v(0)=\delta.$
By Theorem 2.1, there exists a positive entire solution $(u,v)$ of (3.1)–(3.2) with these initial values:
$u(r)$ $=\alpha_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,f(u(s),v(s))\,ds\,dt,$
$v(r)$ $=\delta+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,g(u(s),v(s))\,ds\,dt,$
for $0\leq r<\infty$.
Step 3: Differential inequality for $u+v$. Since $u(r)\geq\alpha_{0}>\delta$ and $v(r)\geq\delta$ for all $r\geq 0$, we have
$u(r)+v(r)\geq\alpha_{0}+\delta>2\delta.$
By assumption (H3), $f(s,t)+g(s,t)\geq h(s+t)$ for $s,t\geq\eta$ (and here $\eta\leq\delta$), so
$\Delta(u+v)=p(r)f(u,v)+q(r)g(u,v)\geq\widetilde{p}(r)\,h(u+v)\quad\text{in }[0,R).$
Step 4: Comparison with $w$ and contradiction. The function $w$ satisfies $\Delta w=\widetilde{p}(r)\,h(w)$ in $[0,R)$, $w(0)=\delta$, and blows up at $r=R$. By the strong maximum principle applied to u+vwu+v-w, if $u(0)+v(0)>w(0)$, then $u+v>w$ near $r=0$. However, since $w$ blows up at $R$ while $u+v$ is $C^{2}$ on $[0,R)$, the comparison principle forces $u+v\leq w$ in $[0,R)$ a contradiction with $u(0)+v(0)>w(0)$.
Therefore, $\mathcal{G}_{\delta}$ cannot be unbounded; it is bounded.
Part II: Closedness of $\mathcal{G}_{\delta}$.
Step 1: Sequential closure argument. Let $\{(\alpha_{m},\beta_{m})\}\subset\mathcal{G}_{\delta}$ be a sequence converging to $(\alpha,\beta)\in\overline{\mathcal{G}_{\delta}}$. For each $m$, Theorem 2.1 guarantees the existence of a positive entire solution $(u_{m},v_{m})$ of (3.1)–(3.2) with
$u_{m}(0)=\alpha_{m},\quad v_{m}(0)=\beta_{m}.$
Step 2: Uniform bounds and equicontinuity. From the monotone iteration scheme in Theorem 2.1 and the Keller–Osserman supersolution $w$ from (3.3) with $w(0)>\sup_{m}(\alpha_{m}+\beta_{m})$, we deduce that
$u_{m}(r)+v_{m}(r)\leq w(r)\quad\forall r\geq 0,\,\forall m\geq 1.$
Hence the family $\{u_{m}\}$ and $\{v_{m}\}$ is uniformly bounded on each compact interval $[0,R]$.
To establish equicontinuity, we use the integral representation (3.1)–(3.2). For $r_{1},r_{2}\in[0,R]$ with $r_{1}<r_{2}$, we have
$|u_{m}(r_{2})-u_{m}(r_{1})|$ $\leq\int_{r_{1}}^{r_{2}}t^{1-n}\int_{0}^{t}s^{n-1}p(s)f(u_{m}(s),v_{m}(s))\,ds\,dt$
$\leq\int_{r_{1}}^{r_{2}}t^{1-n}\int_{0}^{t}s^{n-1}(p(s)+q(s))S(w(s))\,ds\,dt,$
where $S(s)=f(s,s)+g(s,s)$. Since the right-hand side is independent of $m$ and tends to zero as $r_{2}-r_{1}\to 0$ (by absolute continuity of the integral), the family $\{u_{m}\}$ is equicontinuous on $[0,R]$. The same argument applies to $\{v_{m}\}$.
Step 3: Compactness and passage to the limit. By the Arzelà–Ascoli theorem, the sequence $\{(u_{m},v_{m})\}$ is relatively compact in $C([0,R])\times C([0,R])$ for each $R>0$. By a standard diagonal argument, there exists a subsequence $(u_{m_{j}},v_{m_{j}})$ converging uniformly on every compact interval $[0,R]$ to a limit $(u,v)\in C([0,\infty))\times C([0,\infty))$.
To verify that $(u,v)$ satisfies (3.1)–(3.2), fix $r>0$ and observe that the integrands in (3.1) converge:
$s^{n-1}p(s)f(u_{m_{j}}(s),v_{m_{j}}(s))\to s^{n-1}p(s)f(u(s),v(s))$
pointwise in $s\in[0,r]$ by continuity of $f$, and are uniformly bounded by $s^{n-1}(p(s)+q(s))S(w(s))\in L^{1}([0,r])$. By the dominated convergence theorem,
$\lim_{j\to\infty}\int_{0}^{t}s^{n-1}p(s)f(u_{m_{j}}(s),v_{m_{j}}(s))\,ds=\int_{0}^{t}s^{n-1}p(s)f(u(s),v(s))\,ds$
uniformly in $t\in[0,r]$. Applying dominated convergence again to the outer integral in (3.1), we conclude that $(u,v)$ satisfies (3.1)–(3.2), hence is a positive entire solution with
$u(0)=\lim_{j\to\infty}u_{m_{j}}(0)=\lim_{j\to\infty}\alpha_{m_{j}}=\alpha,\quad v(0)=\beta.$
Since $\min\{\alpha,\beta\}\geq\delta$ by construction (as the limit of a sequence in $\mathcal{G}_{\delta}$), we have $(\alpha,\beta)\in\mathcal{G}_{\delta}$.
Conclusion: $\mathcal{G}_{\delta}$ is both bounded and closed, completing the proof of Theorem 2.2. ∎
3.3 Proof of Theorem 2.3
Proof.
We structure the proof into seven distinct and logically self-contained steps.
Standing assumptions and notation. We work under hypotheses (H1)–(H3) and consider the radial system (1.1) in its equivalent integral (Green) form (3.1)–(3.2), with prescribed central values
$(u(0),v(0))=(\alpha,\beta)\in\mathcal{E}(\mathcal{G}_{\delta}).$
We introduce the Keller–Osserman transform
$H(s):=\int_{s}^{\infty}\frac{dt}{f(t,t)+g(t,t)},\quad s>0.$
By construction:
  • $H$ is strictly decreasing, with
    $H^{\prime}(s)=-\frac{1}{f(s,s)+g(s,s)}<0;$
  • $H^{\prime\prime}(s)\geq 0$ since $s\mapsto f(s,s)+g(s,s)$ is nondecreasing by (H2). Explicitly,
    $H^{\prime\prime}(s)=\frac{d}{ds}\left(-\frac{1}{f(s,s)+g(s,s)}\right)=\frac{\frac{\partial}{\partial s}[f(s,s)+g(s,s)]}{[f(s,s)+g(s,s)]^{2}}\geq 0.$
Step 1: Perturbation of initial data and blow-up for approximants. From Theorem 2.1, $\mathcal{G}$ is nonempty and downward closed in each coordinate. From Theorem 2.2, $\mathcal{G}_{\delta}$ is closed and bounded for $\delta>0$ sufficiently small.
Let $(\alpha,\beta)\in\mathcal{E}(\mathcal{G}_{\delta})$. For each $k\geq 1$, consider the perturbed initial data
$\left(\alpha+\frac{1}{k},\ \beta+\frac{1}{k}\right).$
If this pair belonged to $\mathcal{G}$, then by downward closure and the definition of $\mathcal{G}_{\delta}$, $(\alpha,\beta)$ would lie in the interior of $\mathcal{G}_{\delta}$, contradicting $(\alpha,\beta)\in\mathcal{E}(\mathcal{G}_{\delta})$. Hence
$\left(\alpha+\frac{1}{k},\ \beta+\frac{1}{k}\right)\notin\mathcal{G}.$
Therefore, the solution $(u_{k},v_{k})$ of
$u_{k}(r)$ $=\alpha+\frac{1}{k}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)f(u_{k}(s),v_{k}(s))\,ds\,dt,$
$v_{k}(r)$ $=\beta+\frac{1}{k}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)g(u_{k}(s),v_{k}(s))\,ds\,dt,$
must blow up in finite radius: there exists $R_{k}>0$ such that
$\lim_{r\rightarrow R_{k}^{-}}u_{k}(r)=\infty\quad\text{or}\quad\lim_{r\rightarrow R_{k}^{-}}v_{k}(r)=\infty.$
In either case,
$\lim_{r\rightarrow R_{k}^{-}}\big(u_{k}(r)+v_{k}(r)\big)=\infty.$ (3.6)
Moreover, by monotonicity in $k$ of the blow-up radii (see [1]),
$R_{1}\leq R_{2}\leq\dots\leq R_{k}\leq\dots,$
and we set
$R:=\lim_{k\rightarrow\infty}R_{k}\in[R_{1},\infty].$
Step 2: Reduction to a scalar inequality. Let
$W_{k}(r):=u_{k}(r)+v_{k}(r),\quad W(r):=u(r)+v(r),$
where $(u,v)$ is the limit solution corresponding to $(\alpha,\beta)$. Since $u_{k}\leq W_{k}$ and $v_{k}\leq W_{k}$, and $f,g$ are nondecreasing, we have
$\Delta W_{k}$ $=p(r)f(u_{k},v_{k})+q(r)g(u_{k},v_{k})$
$\leq(p(r)+q(r))\,\big(f(W_{k},W_{k})+g(W_{k},W_{k})\big),$ (3.7)
by (H2).
Step 3: Applying the Keller–Osserman transform. Applying $H$ to $W_{k}$ and using $H^{\prime\prime}\geq 0$, the chain rule gives
$\Delta H(W_{k})=H^{\prime}(W_{k})\,\Delta W_{k}+H^{\prime\prime}(W_{k})|\nabla W_{k}|^{2}\geq H^{\prime}(W_{k})\,\Delta W_{k}.$ (3.8)
Justification of the chain rule: Since $W_{k}\in C^{2}$ and $H\in C^{2}((0,\infty))$, the composition $H\circ W_{k}$ is $C^{2}$ wherever $W_{k}>0$. The Laplacian of the composition is
$\Delta H(W_{k})=H^{\prime}(W_{k})\Delta W_{k}+H^{\prime\prime}(W_{k})|\nabla W_{k}|^{2},$
which is the standard chain rule for the Laplacian. Since $H^{\prime\prime}(W_{k})\geq 0$ by our earlier computation, we obtain (3.8).
Since $H^{\prime}(s)=-1/(f(s,s)+g(s,s))$, inequality (3.7) yields
$\Delta H(W_{k})\geq-(p(r)+q(r)).$
Step 4: First radial integration. In radial coordinates, the above inequality becomes
$\big(r^{n-1}(H(W_{k}))^{\prime}(r)\big)^{\prime}\geq-r^{n-1}(p(r)+q(r)).$
Integrating from 0 to $r<R_{k}$ and using regularity at $r=0$ (which implies $\lim_{t\downarrow 0}t^{n-1}(H(W_{k}))^{\prime}(t)=0$), we obtain
$(H(W_{k}))^{\prime}(r)\geq-r^{1-n}\int_{0}^{r}s^{n-1}(p(s)+q(s))\,ds.$
Step 5: Second radial integration and lower bound. Integrating from $r$ to $R_{k}$ gives
$H(W_{k}(R_{k}))-H(W_{k}(r))\geq-\int_{r}^{R_{k}}t^{1-n}\left(\int_{0}^{t}s^{n-1}(p(s)+q(s))\,ds\right)dt.$
Letting $r\rightarrow R_{k}^{-}$, we have
$W_{k}(R_{k})\rightarrow\infty\text{ and hence }H(W_{k}(R_{k}))\rightarrow 0.$
Thus
$H(W_{k}(r))\leq\int_{r}^{R_{k}}t^{1-n}\left(\int_{0}^{t}s^{n-1}(p(s)+q(s))\,ds\right)dt.$
Since $H$ is strictly decreasing, we invert to obtain
$W_{k}(r)\geq H^{-1}\!\left(\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}(p(s)+q(s))\,ds\,dt\right).$
Passing to the limit $k\rightarrow\infty$ yields
$W(r)\geq H^{-1}\!\left(\int_{r}^{R}t^{1-n}\int_{0}^{t}s^{n-1}(p(s)+q(s))\,ds\,dt\right).$
If $R<\infty$, the right-hand side tends to $\infty$ as $r\rightarrow R^{-}$, contradicting the fact that $(u,v)$ is entire. Hence $R=\infty$, and we have established the quantitative lower bound (2.2).
Step 6: Divergence of at least one component. From the previous steps (in particular, Step 1 of the proof of Theorem 2.2 and (3.6)), we already know that
$W(r):=u(r)+v(r)\;\longrightarrow\;+\infty\quad\text{as }r\rightarrow\infty.$
Suppose, for contradiction, that both $u$ and $v$ are bounded on $0,\infty)$. That is, there exist constants $M_{1},M_{2}>0$ such that
$u(r)\leq M_{1},\quad v(r)\leq M_{2}\quad\forall r\geq 0.$
Then
$W(r)\leq M_{1}+M_{2}\quad\forall r\geq 0,$
which contradicts $W(r)\rightarrow+\infty$. Therefore, at least one of the two components must diverge:
$\lim_{r\rightarrow\infty}u(r)=+\infty\quad\text{or}\quad\lim_{r\rightarrow\infty}v(r)=+\infty.$
This next figure shows
$u\left(r\right)\text{ and }v\left(r\right)\text{ vs. }r,$
with the “both bounded” scenario impossible, and the Keller–Osserman case forcing both curves upward.
[…]
Figure 2: $W(r)\rightarrow+\infty$ forces at least one component to diverge; under the Keller–Osserman condition both diverge. Solid lines: without KO; dashed lines: with KO.
Both components diverge under the supplementary Keller–Osserman condition. Assume now that the additional Keller–Osserman condition at infinity (2.3) holds. We claim that in this case $\lim_{r\rightarrow\infty}u(r)\;=\;\lim_{r\rightarrow\infty}v(r)\;=\;+\infty$. Indeed, if $u$ were bounded, then $v=W-u\rightarrow+\infty$. Thus it suffices to prove the implications
$v(r)\rightarrow+\infty\ \Longrightarrow\ u(r)\rightarrow+\infty,\qquad u(r)\rightarrow+\infty\ \Longrightarrow\ v(r)\rightarrow+\infty.$
Claim 1: If $v(r)\rightarrow+\infty$ as $r\rightarrow\infty$, then $u(r)\rightarrow+\infty$ as $r\rightarrow\infty$.
Proof of Claim 1. Assume by contradiction that $u$ is bounded: there exists $U>0$ such that
$0<u(r)\leq U\quad\forall r\geq 0.$
Since $v(r)\rightarrow+\infty$ while $u$ is bounded, there exists $R_{0}>0$ such that
$u(r)\leq U<v(r),\forall r\geq R_{0}.$
By the monotonicity of $f$ in each variable (assumption (H2)), we have
$f\big(u(r),v(r)\big)>f\big(u(r),u(r)\big)=:\ell_{f}\big(u(r)\big)\quad\forall r\geq R_{0},$
where $\ell_{f}(s):=f(s,s)$ is continuous, nondecreasing, satisfies $\ell_{f}(0)=0$, and, by hypothesis, the Keller–Osserman condition
$\int_{1}^{\infty}\frac{dt}{\sqrt{\int_{0}^{t}\ell_{f}(z)\,dz}}<\infty.$
Thus, for $r\geq R_{0}$,
$\Delta u(r)>p(r)\,\ell_{f}\big(u(r)\big).$ (3.9)
Barrier construction. Fix $R_{0}>0$ as above and consider the radial Cauchy problem
$y^{\prime\prime}(r)+\frac{n-1}{r}y^{\prime}(r)$ $=p(r)\,\ell_{f}\big(y(r)\big),\quad r\in[R_{0},R_{a}),$ (3.10)
$y(R_{0})$ $=a<u(R_{0}),\quad y^{\prime}(R_{0})=0,$
which has a maximal $C^{2}$ solution $y$ on $R_{0},R_{a})$ by standard ODE theory.
Define the Keller–Osserman primitive
$\mathcal{C}_{f}(s):=\int_{s_{\ast}}^{s}\frac{d\tau}{\ell_{f}(\tau)},\quad s>s_{\ast},$
where $s_{\ast}>0$ is chosen so that
$\ell_{f}(\tau)>0\text{ on }[s_{\ast},\infty).$
Then $\mathcal{C}_{f}$ is strictly increasing on $(s_{\ast},\infty)$, with
$\mathcal{C}_{f}^{\prime}(s)=\frac{1}{\ell_{f}(s)},\quad\mathcal{C}_{f}^{\prime\prime}(s)=-\frac{\ell_{f}^{\prime}(s)}{\ell_{f}(s)^{2}}\leq 0,$
and, by the Keller–Osserman condition, $\mathcal{C}_{f}(\infty)<\infty$.
Set $W(r):=\mathcal{C}_{f}\big(y(r)\big)$. Differentiating and using (3.10) yields
$W^{\prime}(r)=\frac{y^{\prime}(r)}{\ell_{f}(y(r))},\quad\big(r^{n-1}W^{\prime}(r)\big)^{\prime}=\frac{r^{n-1}}{\ell_{f}(y)}\left(y^{\prime\prime}+\frac{n-1}{r}y^{\prime}\right)-\frac{r^{n-1}\ell_{f}^{\prime}(y)}{\ell_{f}(y)^{2}}\,(y^{\prime})^{2}\leq r^{n-1}p(r).$
Integrating from $R_{0}$ to $r$ and using $W^{\prime}(R_{0})=y^{\prime}\left(R_{0}\right)/\mathcal{C}_{f}(a)=0$ gives
$r^{n-1}W^{\prime}(r)\leq\int_{R_{0}}^{r}s^{n-1}p(s)\,ds.$
Dividing by $r^{n-1}$ and integrating again from $R_{0}$ to $r$ yields the Keller–Osserman estimate
$\mathcal{C}_{f}\big(y(r)\big)\leq\mathcal{C}_{f}(a)+Q_{R_{0}}(r),\quad Q_{R_{0}}(r):=\int_{R_{0}}^{r}t^{1-n}\left(\int_{R_{0}}^{t}s^{n-1}p(s)\,ds\right)dt.$ (3.11)
By direct computation (Fubini’s theorem and $n\geq 3$),
$0\leq Q_{R_{0}}(\infty)=\frac{1}{n-2}\int_{R_{0}}^{\infty}\left(1-\left(\frac{R_{0}}{t}\right)^{n-2}\right)t\,p(t)\,dt\leq\frac{1}{n-2}\int_{R_{0}}^{\infty}t\,p(t)\,dt,$
so $Q_{R_{0}}(\infty)<\infty$ and $Q_{R_{0}}(\infty)\downarrow 0$ as $R_{0}\uparrow\infty$.
Blow-up of the barrier. Since $\mathcal{C}_{f}(\infty)<\infty$, we can first choose $R_{0}$ large so that $Q_{R_{0}}(\infty)$ is as small as desired, then choose $a$ close enough to $u(R_{0})$ but still smaller than $u(R_{0})$, so that $\mathcal{C}_{f}(a)\;+\;Q_{R_{0}}(\infty)\;>\;\mathcal{C}_{f}(\infty)$. Because $Q_{R_{0}}$ is continuous and strictly increasing, there exists a finite $R_{a}\in(R_{0},\infty)$ with
$\mathcal{C}_{f}(a)\;+\;Q_{R_{0}}(R_{a})\;=\;\mathcal{C}_{f}(\infty).$ (3.12)
By (3.11), this forces $\mathcal{C}_{f}\bigl(y(r)\bigr)\nearrow\mathcal{C}_{f}(\infty)$ as $r\uparrow R_{a}$, hence $\lim_{r\uparrow R_{a}}y(r)\;=\;+\infty$, i.e., $y$ is a (finite-radius) blow-up supersolution for (3.9) on $R_{0},R_{a})$.
Maximum principle comparison and contradiction. Fix any $R\in(R_{0},R_{a})$. Since $u$ is assumed bounded on $R_{0},\infty)$ while $y(r)\rightarrow\infty$ as $r\uparrow R_{a}$, we may choose $R$ sufficiently close to $R_{a}$ so that
$y(R)>\sup_{r\geq R_{0}}u(r).$
In particular,
$u(R_{0})-y(R_{0})=u(R_{0})-a>0,\qquad u(R)-y(R)<0.$
Define w:=uyw:=u-y on the annulus
$A_{R_{0},R}:=\bigl\{x\in\mathbb{R}^{n}:\;R_{0}<r=|x|<R\bigr\}\,\}.$
Using (3.9) and (3.10), together with the monotonicity of $l_{f}$, we obtain on the set where $u\geq y$:
$\Delta w=\Delta u-\Delta y>p(r)\,l_{f}(u)-p(r)\,l_{f}(y)\geq 0.$
Let
$r_{\ast}:=\inf\{\,r\in(R_{0},R]:w(r)\leq 0\,\}.$
By the boundary values above, $r_{\ast}\in(R_{0},R)$ and
$w(r)>0\quad\text{for }r\in[R_{0},r_{\ast}),\qquad w(r_{\ast})=0.$
Thus $w$ is nonnegative on $R_{0},r_{\ast}]$ and strictly positive inside, with $w$ attaining its minimum 0 at the boundary point $r_{\ast}$ of the domain $(R_{0},r_{\ast})$. Moreover, on $(R_{0},r_{\ast})$ we have $u\geq y$ and hence $\Delta w>0$.
The situation is illustrated in Figure, where the blue curve denotes $u$, the red curve the barrier $y$.
[…]
Figure 3: Schematic of the first contact $r_{\ast}$ where $u$ (blue) and $y$ (red) intersect: solid for $r<r_{\ast}$, dashed for $r>r_{\ast}$. On $(R_{0},r_{\ast})$ we have w=uy>0w=u-y>0 and $\Delta w>0$, so the Hopf boundary lemma yields $w^{\prime}(r_{\ast})>0$, contradicting the first-contact requirement $w^{\prime}(r_{\ast})\leq 0$.
By the (radial) weak maximum principle applied to $w$ on $(R_{0},r_{\ast})$, $w$ cannot achieve a strict interior minimum; by the Hopf boundary point lemma at $r_{\ast}$, the outward normal derivative must satisfy
$w^{\prime}(r_{\ast})>0\text{ (outward normal points to ward increasing }r\text{)}.$
However, $r_{\ast}$ is the first contact point where $w$ crosses from positive to nonpositive values, so necessarily $w^{\prime}(r_{\ast})\leq 0$. This contradiction shows that such $r_{\ast}$ cannot exist. The next Figure shows $u$ vs. $y$ and why the “forbidden crossing” at $r_{\ast}$ can’t happen.
[…]
Figure 4: $u=y$ separates the regions $u>y$ and $u<y$. The trajectory starts in $u>y$ at $R_{0}$ and touches the diagonal for the first time at $r_{\ast}$. The maximum principle forbids crossing into $u<y$ here.
Hence w(r)=u(r)y(r)>0w(r)=u(r)-y(r)>0 on $R_{0},R)$. Letting $R\uparrow R_{a}$ yields $u(r)\geq y(r)$ for all $r\in[R_{0},R_{a})$. Since $y(r)\rightarrow\infty$ as $r\uparrow R_{a}$, we must have $u(r)\rightarrow\infty$ as $r\uparrow R_{a}$, which contradicts the assumed boundedness of $u$ (and contradicts the prescribed initial value comparison $u(R_{0})>a$ together with the smoothness of $u$ on $R_{0},\infty)$). Therefore, the assumption that $u$ is bounded while $v\rightarrow\infty$ is impossible.
Claim 2: If $u(r)\rightarrow+\infty$ as $r\rightarrow\infty$, then $v(r)\rightarrow+\infty$ as $r\rightarrow\infty$.
Proof of Claim 2. The argument is symmetric, using $\ell_{g}$, $q$, and a barrier $w$ with $w(R_{1})<v(R_{1})$ to obtain the same type of contradiction. ∎
Step 7: Conclusion. We have now completed the construction of a positive, radial, entire solution
$(u,v)\in C^{2}([0,\infty))\times C^{2}([0,\infty))$
to the system (1.1) under the standing assumptions (H1)–(H3). From the previous steps, we have established that the sum
$W(r):=u(r)+v(r)$
satisfies
$W(r)\longrightarrow+\infty\quad\text{as }r\rightarrow\infty,$
and moreover obeys the quantitative lower bound (2.2) derived via the Keller–Osserman transform.
We have also shown that:
  • Without any additional conditions added to (H1)-(H3), at least one of the two components $u$ or $v$ must diverge to $\infty$ as $r\rightarrow\infty$.
  • Under the supplementary Keller–Osserman condition at infinity (2.3), both components necessarily diverge:
    $\lim_{r\rightarrow\infty}u(r)=\lim_{r\rightarrow\infty}v(r)=+\infty.$
Thus, the constructed pair $(u,v)$ is a positive, radial, entire solution of (1.1) with the stated asymptotic properties, completing the proof of Theorem 2.3. ∎
4 Conclusion
In this work, we have extended the existence theory for entire large solutions of competitive semilinear elliptic systems beyond the classical power-law setting studied by Lair [3]. Our generalization applies to a broad class of nonlinearities that satisfy a Keller–Osserman-type growth condition, thereby encompassing both critical and supercritical regimes within a unified analytical framework.
The proof strategy preserves the essential structure of Lair’s original approach, while adapting each step to handle the greater generality of the nonlinear terms:
  • We employ a monotone iteration scheme to construct a global positive radial solution starting from admissible central values.
  • By summing the two equations, we reduce the system to a scalar differential inequality for $W=u+v$, which captures the combined growth of both components.
  • We apply a Keller–Osserman transform to $W$, exploiting the monotonicity and convexity properties of the transform to control the solution’s behaviour.
  • A two-step radial integration argument then forces blow-up of $W$ at infinity, ensuring that at least one component or both of the solution diverges.
This approach not only recovers the known results for pure power nonlinearities as a special case, but also provides a flexible method applicable to more general growth laws. In particular, the Keller–Osserman framework allows us to treat critical and supercritical growth in a single, coherent setting, highlighting the structural mechanisms that drive the existence of large solutions in competitive elliptic systems.
Future directions. Several natural extensions of this work merit investigation:
  • Asymptotic profiles: While we establish lower bounds on the blow-up rate, determining the precise asymptotic behavior of $(u,v)$ as $r\to\infty$ remains open for general nonlinearities.
  • Non-radial solutions: Our methods are inherently radial. Extending to non-radial entire solutions would require fundamentally different techniques.
  • Refined Keller–Osserman conditions: Investigating the sharpness of condition (H3) and exploring intermediate growth regimes could yield interesting borderline phenomena.
  • Systems with more than two equations: Generalizing to $m$-component systems ($m\geq 3$) presents both technical and conceptual challenges.
Acknowledgments
The author thanks the anonymous referees for their careful reading and helpful suggestions that improved the presentation of this paper.